### v1

# Spectroscopy and complex-time correlations using minimally entangled typical thermal states

Abstract
Tensor network states have enjoyed great success at capturing aspects of strong correlation physics. However, obtaining dynamical correlators at non-zero temperatures is generically hard even using these methods. Here, we introduce a practical approach to computing such correlators using minimally entangled typical thermal states (METTS). While our primary method directly computes dynamical correlators of physical operators in real time, we propose extensions where correlations are evaluated in the complex-time plane. The imaginary time component bounds the rate of entanglement growth and strongly alleviates the computational difficulty allowing the study of larger system sizes. To extract the physical correlator one must take the limit of purely real-time evolution. We present two routes to obtaining this information (i) via an analytic correlation function in complex time combined with a stochastic analytic continuation method to obtain the real-time limit and (ii) a hermitian correlation function that asymptotically captures the desired correlation function quantitatively without requiring effort of numerical analytic continuation. We show that these numerical techniques capture the finite-temperature dynamics of the Shastry-Sutherland model -- a model of interacting spin one-half in two dimensions.
I Introduction
Understanding the macroscopic behavior of quantum materials from microscopic descriptions of strongly interacting particles is one of the great challenges of contemporary physics. Developments in numerical simulations have given us powerful tools to gain insights beyond systems where an analytical treatment is still feasible. While several theoretical models in condensed matter physics are amenable to highly accurate quantum Monte Carlo (QMC) simulations [1], others like frustrated magnets or generic strongly interacting fermion systems present us with the sign problem when tackled using QMC [2], limiting the range of parameters and observables that can be studied. Tensor network methods, on the other hand, are not afflicted by a sign problem [3, 4]. Instead, these methods face difficulties whenever confronted with the problem of dealing with highly entangled states. The arguably most well-known tensor network method is the density matrix renormalization group (DMRG) [5, 6, 7] which, in its modern version, is based on matrix product states (MPS) [8]. Initially proposed as a method to study static properties of ground states, several extensions thereof have enabled the study of static observables at finite temperature [9, 10, 11, 12, 13, 14, 15, 16, 17, 18], dynamical spectral functions at zero temperature [19, 20, 21, 22, 23, 24, 25, 26, 27, 28], and unitary dynamics in non-equilibrium settings [29, 21, 30, 31, 32, 33, 34, 35].
To fully build a bridge between microscopic models and their experimental response functions, simulations of dynamical correlators at finite temperature are necessary. This includes dynamical spin structure factors as probed in neutron scattering experiments [36, 37], and electronic Green’s functions probed in angular resolved photoemission spectroscopy (ARPES) [38]. Moreover, transport coefficients can be understood within the Kubo linear response formalism. The computation of dynamical response functions at finite temperature in two spatial dimensions, however, still poses significant difficulties for tensor network methods.
A well-known DMRG/MPS-based approach to compute time-dependent correlation functions is to perform real-time evolution. This technique has been applied at zero temperature [21, 26, 27], but also approaches at finite temperature using purifications have been reported [39, 40, 41, 42, 43, 44] as well as an interesting approach to investigate high-temperature transport using dissipation-assisted operator evolution [45]. For alternative approaches to finite-$T$ dynamics, see also Refs. [46, 47].
The real-time evolution approach, however, suffers from the fact that the entanglement of generic quantum systems grows unbounded if evolved for long times. In practice, this limits the system size which can be simulated as well as the resolution of the spectral function which is attained. The computation of real-time correlators is also especially tricky in QMC simulations, where an even more intricate phase problem is encountered. Interestingly, however, much of many-body theory is built on the notion of imaginary-time correlators. While these correlations do not immediately allow for the extraction of dynamical response functions to arbitrary precision at any real frequency, properties like excitation gaps can be reliably estimated [48]. Within the framework of MPS techniques, the computation of imaginary-time evolution does not face the same problem of unbounded entanglement growth. In fact, at asymptotically long times imaginary time evolution (under mild assumptions) yields the many-body ground state, whose entanglement for most physically relevant systems is expected to be small as compared to genuine quantum states with volume-law entanglement. Famously, ground states of locally interacting and gapped systems obey an area law [49, 50]. Thus, imaginary-time evolution poses fewer challenges as compared to real-time evolution when applying MPS techniques, which allows one to study larger systems and a wider range of parameters.
In this paper, we introduce an approach to evaluating dynamical spectral functions based on the minimal entangled typical thermal states (METTS) [12, 13, 14, 15, 16, 17] technique to simulate systems at nonzero temperature. Moreover, we propose to use a mixed approach where correlators are evaluated in fully complex time, i.e., having both a real-time and imaginary-time component. We present efficient algorithms to obtain the dynamical spectrum of physical operators at nonzero temperature, by computing correlation functions along a certain angle in the complex time plane. Such an approach has originally been proposed in Ref. [51] for QMC simulations but has until now not been widely adopted. Recently, however, complex time correlation functions have been revisited in the context of MPS ground state simulations and applied to quantum impurity models [52, 53].
As we will show, complex-time correlation functions can be evaluated using well-known MPS time evolution algorithms such as the time-dependent variation principle (TDVP) [54, 55]. Our aim is to combine the advantages of computing imaginary-time correlations and real-time correlations. The imaginary contribution of time evolution dissipates the large entanglement entropy while the real-time part of the correlation function benefits the efficiency of obtaining real frequency spectral function.
We compare two approaches that generalize time evolution to complex times. The first approach generalizes the imaginary time correlator that is well-discussed in the QMC community to the complex plane in an analytic way, such that the spectral functions in real frequency can be obtained following a numerical analytical continuation method. In the second approach, we introduce a hermitian complex correlation function, which can be implemented analogously to the real-time correlation function.
[…]
The outline of our paper is as follows. We begin with Sec. II introducing the Shastry-Sutherland model we use to benchmark out methods. In Sec. III.1, we review the conventional thermal METTS algorithm of simulating static observables at nonzero temperatures. The dynamical METTS algorithm of obtaining spectral function via real-time correlations is introduced in Sec. III.2. As both these algorithms are based on efficient matrix-product state time-evolution algorithms, we discuss the time-evolution methods employed in Sec. III.3. Details on the Fourier analysis of time-dependent correlation functions are discussed in Sec. III.4. To validate our approach, we perform a comparison of the dynamical METTS algorithm to spectral functions obtained via the Lehmann representation using full exact diagonalization and DMRG in Sec. III.5.
In the second part of our publication, we discuss how the computational cost from real-time evolution can be eased by performing genuinely complex-time evolution. In Sec. IV.1, we introduce two definitions of complex time correlation functions. One of these, that we call the analytic correlation function is an interpolation between the real-time correlation function and the imaginary-time Matsubara Green’s functions often employed in many-body physics. We discuss an efficient algorithm evaluating this correlation function using METTS in Sec. IV.2. From these correlation functions, the real-time spectral function can be obtained via analytic continuation. We describe our approach employing stochastic analytic continuation in Sec. IV.3. A second interesting kind of complex time correlation, we call the hermitian correlation function is introduced in Sec. IV.4. We show that this correlation can be efficiently evaluated with bounded entanglement while being a good approximation to the real-time correlation function. Finally, we give a summary and perspective in Sec. V.
[…]
The Shastry-Sutherland model [57, 58] is chosen to showcase and benchmark our numerical method. The model is given by,
$H=J_{D}\sum_{\langle i,j\rangle}\bm{S}_{i}\cdot\bm{S}_{j}+J\sum_{\langle\langle i,j\rangle\rangle}\bm{S}_{i}\cdot\bm{S}_{j}.$ (1)
where $\bm{S}_{i}=(S^{x}_{i},S^{y}_{i},S^{z}_{i})$ denotes the spin-$1/2$ operators, $J_{D}$ denotes the intradimer coupling and $J$ denotes the interdimer coupling on the Shastry-Sutherland lattice, see Fig. 1.
[…]
We consider cylindrical geometries of length $L$ and width $W$ with periodic boundary conditions in the short $y$-direction and open boundary conditions in the longer $x$-direction. In this work, we focus on width $W=4$ and $W=6$ cylinders with length $L=16$. A peculiarity when using open boundary conditions for the Shastry-Sutherland model is that there exist spins that are not coupled by any intradimer coupling $J_{D}$ and only with interdimer couplings $J$. Flipping such a dangling spin would constitute an artificial boundary mode and, therefore, we remove interdimer couplings that couple to these boundary dangling spins.
Figure 2: Illustration of the dynamical METTS algorithm. A sequence of product states $|\sigma_{i}\rangle$ and METTS $|\psi_{i}\rangle$ is sampled according to the thermal METTS algorithm. For every METTS the time-dependent correlation function $\mathcal{C}_{AB}^{i}(t)$ as in Eq. 20 is computed. Product states $|\sigma_{i+1}\rangle$ are sampled with transition probability $T_{i\rightarrow i+1}=|\langle\psi_{i}|\sigma_{i+1}\rangle|$. From the individual METTS correlation functions $\mathcal{C}_{AB}^{i}(t)$ the spectral function $\mathcal{C}_{AB}(\omega)$ is computed via a Fourier transform and subsequent averaging.
We are interested in the dynamical spin structure factor given by
[…]
where $S^{z}(\bm{q},t)=e^{iHt}S^{z}(\bm{q})e^{-iHt}$ and,
$S^{z}(\bm{q})=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}e^{i\bm{q}\cdot\bm{r}_{i}}S^{z}_{i},$ (3)
where $\bm{r}_{i}$ denotes the position of site $i$ and $N$ denotes the number of lattice sites. The position of the lattice sites $\bm{r}_{i}$ as well as the unit cell and basis sites are considered in the square lattice geometry shown in Fig. 1(b). To reduce the boundary effects of the cylinder we use a modified definition,
$\tilde{S}^{z}(\bm{q})=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}w(\bm{r}_{i})e^{i\bm{q}\cdot\bm{r}_{i}}S^{z}_{i},$ (4)
where $w(\bm{r})=w(x,y)$ denotes a windowing function. We choose a Tukey window of width $a=4$ as defined by,
$w(x,y)=$ (5)
$\begin{cases}\frac{1}{2}\left\{1+\cos\left(\frac{\pi}{a}(x-a)\right)\right\}&\text{ if }0\leq x\leq a\\ \frac{1}{2}\left\{1+\cos\left(\frac{\pi}{a}(x-(L-1)+a)\right)\right\}&\text{ if }x\geq(L-1)-a\\ 1&\text{ otherwise. }\\ \end{cases}$
Explicitly, for $L=16$ and $a=4$ this means $w(0,y)=w(L-1,y)=0$, $w(1,y)=w(L-2,y)\approx 0.14644$, $w(2,y)=w(L-3,y)=0.5$, $w(3,y)=w(L-4,y)\approx 0.85355$, and $w(x,y)=1$ else. For notational simplicity, $S(\bm{q},\omega)$ will, in the following, refer to the spectral function employing the windowed definition of $\tilde{S}^{z}(\bm{q})$ instead of $S^{z}(\bm{q})$, whenever cylindrical boundary conditions are used.
III Dynamical METTS algorithm
After first reviewing the thermal METTS algorithm we introduce the dynamical METTS algorithm. We discuss computational details such as the appropriate choice of MPS time evolution algorithms and windowing for Fourier transformation of data simulated up to a finite time. The algorithm is then validated against results from exact diagonalization and DMRG.
III.1 Review of the thermal METTS algorithm
The METTS algorithm [12, 13] computes thermal expectation values of any operator $\mathcal{O}$ by sampling over expectation values of certain pure states. Generically, we can write a thermal expectation value as,
$\langle\mathcal{O}\rangle_{\beta}$ $=\frac{1}{\mathcal{Z}}\text{Tr}[e^{-\beta H}\mathcal{O}]$ (6)
$=\frac{1}{\mathcal{Z}}\sum_{i}\langle\sigma_{i}|e^{-\beta H/2}\mathcal{O}e^{-\beta H/2}|\sigma_{i}\rangle$ (7)
$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\langle\psi_{i}|\mathcal{O}|\psi_{i}\rangle,$ (8)
where $\beta=1/(k_{\text{B}}T)$ (henceforth $k_{\text{B}}=1$) denotes the inverse temperature, $\mathcal{Z}=\text{Tr}[e^{-\beta H}]$ the partition function, and $\braket{\ldots}_{\beta}=\frac{1}{\mathcal{Z}}\text{Tr}\left[e^{-\beta H}\ldots\right]$ denotes a thermal expectation value. The sum is performed over the computational basis of product states,
$|\sigma_{i}\rangle=|\sigma_{i}^{1}\rangle|\sigma_{i}^{2}\rangle\;\ldots\;|\sigma_{i}^{N}\rangle,$ (9)
and we define the METTS state,
$|\psi_{i}\rangle=\frac{1}{\sqrt{p_{i}}}e^{-\beta H/2}|\sigma_{i}\rangle,$ (10)
where $p_{i}=\langle\sigma_{i}|e^{-\beta H}|\sigma_{i}\rangle\geq 0$ is is a (non-negative) weight. Notice that the partition function can also be written as $\mathcal{Z}=\sum_{i}p_{i}$. The METTS states are normalized, i.e., $|\langle\psi_{i}|\psi_{i}\rangle|^{2}=1$. To sample the METTS states $|\psi_{i}\rangle$ with probability $p_{i}/\mathcal{Z}$ a Markov chain is constructed where the transition probability is given by,
$T_{i\rightarrow j}=|\langle\psi_{i}|\sigma_{j}\rangle|^{2}.$ (11)
The original works on METTS [12, 13] proved that this transition probability fulfills the detailed balance conditions,
$p_{i}T_{i\rightarrow j}=p_{j}T_{j\rightarrow i},$ (12)
and, therefore, the stationary distribution of this Markov chain is given by the probabilities $p_{i}$. As a Markov chain Monte Carlo simulation, thermalization and autocorrelation effects have to be taken into account to estimate averages and standard errors. The METTS algorithm is particularly well-suited to tensor network simulations describing states of low entanglement. Due to the fact that we sample with non-negative probabilities $p_{i}$ the Monte Carlo sampling itself is not accompanied by a sign problem, as often encountered in QMC simulations. The computational complexity is rather transferred to the entanglement of the METTS states |$\ket{\psi_{i}}$.
III.2 Dynamical METTS algorithm in real-time
Figure 3: Exemplary METTS samples (blue lines) of (a) the correlation functions $\mathcal{C}_{AB}^{i}(t)$ and (b) the spectral functions $\mathcal{C}_{AB}^{i}(\omega)$ of the Shastry-Sutherland model on a $16\times 4$ cylinder for $J/J_{D}=0.63$ and $T/J_{D}=0.1$. We investigate the spin correlations and structure factors with $A^{\dagger}=B=S^{z}(\bm{q})$ as in Eqs. 3 and 4 with $\bm{q}=(\pi,0)$. The estimated averages $\mathcal{C}_{AB}(t)$ and $\mathcal{C}_{AB}(\omega)$ with a statistical error estimate from $R=100$ samples are shown as the red line with the shaded light red region.
To study dynamical structure factors we are interested in the conventional time-dependent correlation function $\mathcal{C}_{AB}(t)$ given by,
$\mathcal{C}_{AB}(t)=\langle A(t)B\rangle_{\beta}=\langle e^{iHt}Ae^{-iHt}B\rangle_{\beta},$ (13)
where $A$ and $B$ denote observables subject to investigation. For the dynamical spin structure factor considered in this work, we choose,
$A^{\dagger}=B=S^{z}(\bm{q}),$ (14)
as defined in Eqs. 3 and 4. A dynamical spectral function is then given as the Fourier transform,
[…]
We expand the corresponding expectation values in the basis of METTS states,
$\mathcal{C}_{AB}(t)=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\langle\psi_{i}|e^{iHt}Ae^{-iHt}B|\psi_{i}\rangle,$ (16)
and define the states,
$|v_{i}(t)\rangle=e^{-iHt}B|\psi_{i}\rangle,$ (17)
$|w_{i}(t)\rangle=e^{-iHt}|\psi_{i}\rangle.$ (18)
Notice, while the state |$\ket{w_{i}(t)}$ is normalized, we generically have,
$\braket{v_{i}(t)}{v_{i}(t)}=\text{constant}\neq 1.$ (19)
We introduce the time-dependent correlation function of an individual METTS state as,
$\mathcal{C}_{AB}^{i}(t)$ $\equiv\langle\psi_{i}|e^{iHt}Ae^{-iHt}B|\psi_{i}\rangle$ (20)
$=\langle w_{i}(t)|A|v_{i}(t)\rangle.$ (21)
Hence, the time-dependent correlation function and the dynamical spectral function are expressed as,
$\mathcal{C}_{AB}(t)=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\mathcal{C}_{AB}^{i}(t),$ (22)
$\quad\mathcal{C}_{AB}(\omega)=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\mathcal{C}_{AB}^{i}(\omega),$ (23)
where,
[…]
This suggests the following algorithm to compute dynamical correlation functions $\mathcal{C}_{AB}^{i}(t)$ and $\mathcal{C}_{AB}^{i}(\omega)$. We sample METTS states |$\ket{\psi_{i}}$ by sampling from the Markov chain with transition probability $T_{i\rightarrow j}=|\langle\psi_{i}|\sigma_{j}\rangle|^{2}$, just as in the thermal METTS algorithm. For every METTS state |$\ket{\psi_{i}}$ we compute $|v_{i}(0)\rangle=B|\psi_{i}\rangle$ and set $|w_{i}(0)\rangle=|\psi_{i}\rangle$ at the beginning.
Choose a random initial product state $|\sigma_{1}\rangle=|\sigma_{1}^{1}\rangle\cdot\ldots|\sigma_{1}^{N}\rangle$
for $i=1,\ldots,R$ do
     Compute the METTS state $|\psi_{i}\rangle=\frac{1}{\sqrt{p_{i}}}e^{-\beta H/2}|\sigma_{i}\rangle$
     Compute $|v_{i}(0)\rangle=B|\psi_{i}\rangle$ and set $|w_{i}(0)\rangle=|\psi_{i}\rangle$
     Set $t=0$
     while $t\leq\Omega$ do
         Compute $\mathcal{C}^{i}(t)=\langle w_{i}(t)|A|v_{i}(t)\rangle$
         Compute $|v_{i}(t+\delta t)\rangle=e^{-iH\delta t}|v_{i}(t)\rangle$
         Compute $|w_{i}(t+\delta t)\rangle=e^{-iH\delta t}|w_{i}(t)\rangle$
         $t\rightarrow t+\delta t$
     end while
     Store correlator $\mathcal{C}^{i}(t)$ to disk
     Pick next product state $|\sigma_{i+1}\rangle$ with probability $|\langle\psi_{i}|\sigma_{i+1}\rangle|^{2}$
end for
Algorithm 1 Dynamical METTS for correlators $\mathcal{C}(t)$
The next step is to perform a real-time evolution to compute $|v_{i}(t)\rangle$ and $|w_{i}(t)\rangle$ as in Eqs. 17 and 18. Notice that the METTS states $|\psi_{i}\rangle$ are not eigenstates of the Hamiltonian $H$, such that the time-evolved states $|v_{i}(t)\rangle$ and $|w_{i}(t)\rangle$ have to be explicitly computed. Naturally, this time evolution can only be carried out until a maximum time $\Omega$. Measurements of $\mathcal{C}_{AB}^{i}(t)$ are only performed at discrete time steps which are a multiple of an elementary step size $\delta_{t}$.
For convenience, we introduce the shorthand notations,
$C(t)=C_{AB}(t),\quad$ $C(\omega)=C_{AB}(\omega),$ (25)
$C^{i}(t)=C^{i}_{AB}(t),\quad$ $C^{i}(\omega)=C^{i}_{AB}(\omega).$ (26)
The algorithm laid out above is summarized in Alg. 1, and illustrated graphically in Fig. 2. Using the Shastry-Sutherland model as a concrete example, we show the corresponding correlation function $\mathcal{C}_{AB}^{i}(t)$ and $\mathcal{C}_{AB}^{i}(\omega)$ for several sampled METTS configurations as well as their average value in Fig. 3.
The real-time approach of measuring spectral functions at finite temperature is rigorous as in the limit of long time domain and sufficient sampling of METTS states |$\ket{\psi_{i}}$. However, a limitation of this algorithm is naturally the unbounded growth of entanglement entropy in real-time evolution. In particular, at higher temperatures entanglement growth can hinder accurate calculations at given computational resources. As shown in Fig. 6 below, the bipartite entanglement entropy $S_{\text{vN}}$ of time evolution based on METTS states of the Shastry-Sutherland model at $T/J_{D}=0.2$ grows much more quickly on average than the one at $T/J_{D}=0.1$. Meanwhile, pronounced fluctuations of the entropy after time evolution exist among various sampled METTS configurations.
In the limit of infinitely long simulation time $\Omega\rightarrow\infty$ and infinitely accurate calculations of the entangled states |$\ket{\psi_{i}}$, |$\ket{v_{i}(t)}$, and |$\ket{w_{i}(t)}$, this algorithm is exact assuming ergodicity of the METTS Markov chain. A finite time horizon $\Omega$ introduces an artificial broadening to the spectral function on the order of $\propto 1/\Omega$. The accuracy in computing time-evolutions of MPS states can be well controlled by modern algorithms described in the next section.
III.3 Time-evolution using matrix product states
The key computational bottleneck in both the thermal as well as the dynamical METTS algorithm is the computation of either the imaginary-time evolution of the product states |$\ket{\sigma_{i}}$ or the real-time evolutions |$\ket{v_{i}(t)}$ and |$\ket{w_{i}(t)}$ of the METTS state |$\ket{\psi_{i}}$. Research on time-evolution techniques for matrix product states has received a considerable amount of attention over the years resulting in several practical methods that are in common use. In a recent review [67], these methods have been thoroughly compared in various test scenarios. These thorough comparisons suggest that the time-dependent variational principle (TDVP) for MPS [54, 55] often performs optimally, even though other approaches like the WII method [68] can perform equally well.
Similarly to DMRG, the TDVP method can be performed in a single-site (1TDVP) and a two-site version (2TDVP). While the 1TDVP method conserves energy exactly when performing a real-time evolution and is faster than 2TDVP, it does not allow the bond dimension to be increased without further extensions. 2TDVP on the other hand allows the MPS bond dimension to increase during the evolution. Both TDVP methods come with two control parameters. The first one is the time step $\tau$ which can be chosen to be remarkably large while only incurring a small error. In Ref. [14] a detailed study was performed of the accuracy of TDVP depending on the time step size $\tau$. There, a time step size between $\tau=0.1,\ldots,0.2$ in units of the hopping amplitude $t$ of the Hubbard model was found to be optimal. The second control parameter is the cutoff $\varepsilon$ which is the analog of the DMRG truncation error which, here, is chosen to be $10^{-6}$. Interestingly, TDVP does not become more accurate in the limit $\tau\rightarrow 0$, as then more truncations are necessarily leading to an increased overall error.
When performing long real-time evolutions we use 2TDVP to increase the MPS bond dimension until a maximal bond dimension $\chi_{\max}$ is encountered. Thereafter, 1TDVP is employed to further time-evolve the MPS. Results are then studied as a function of $\chi_{\max}$ to investigate convergence. A peculiarity of the TDVP algorithm is that upon evolving states with small bond dimensions (e.g. product states), a so-called projection error is encountered [67]. To avoid this error, the initial time evolution for the thermal METTS states should be performed by a different method with high accuracy. For this purpose, the time-evolving block decimation [69] algorithm has been applied previously [14]. However, more recently a basis extension technique for 1TDVP has been proposed [70], which can be applied easily without necessitating a Trotterization with possible swap gates. In our specific example, at very low temperatures, the projection error leads to a potential issue when evaluating the analytic correlation function away from the real-time axis. This is due to the low entanglement property of the ground state of the Shastry-Sutherland model, and we document this discussion in Appendix A. This issue is specific to the dimer product ground state of the Shastry-Sutherland model and is not expected to occur in general for more entangled ground states. Our implementation is based on the ITensor library [71] and the TDVP package of ITensor [72].
Figure 4: Comparison of results from dynamical METTS to the spectral function from exact diagonalization on a $4\times 4$ lattice with fully periodic boundary conditions at $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$ for temperatures $T/J_{D}=0.1$ (a) and $T/J_{D}=0.2$ (b). As the exact spectral function from ED is a sum of delta functions, we show Gaussian broadenings for widths $\eta=0.02,0.05,0.10$. For the dynamical METTS, we used $R=100$ samples with a real-time evolution up to $\Omega/J_{D}=50$. We observe close agreement between the METTS results and the ED results at Gaussian broadening $\eta=0.05$.
III.4 Spectral analysis and windowing
Evaluating the METTS spectral functions $\mathcal{C}^{i}(\omega)$ via the Fourier transform Eq. 24 cannot be done exactly as we are only provided with data for $\mathcal{C}^{i}(t)$ on discrete times,
[…]
In our simulations, we found that a thorough treatment of the Fourier transform is crucial to arrive at accurate spectral functions. A naive discrete Fourier transform of the form,
$\mathcal{C}^{i}(\omega)\approx\delta_{t}\sum_{k=-S}^{S}e^{-i\omega k\delta_{t}}\mathcal{C}^{i}(k\delta_{t}),$ (28)
leads to artificial spectral content due to the abrupt end of the signal and cannot be used for accurate measurements of the spectral function. Instead, we employ a windowing function $w(k)$ to consider the windowed Fourier transform [73],
$\mathcal{C}^{i}(\omega)\approx\delta_{t}\sum_{k=-S}^{S}e^{-i\omega k\delta_{t}}\mathcal{C}^{i}(k\delta_{t})w(k).$ (29)
In the context of DMRG simulations of spectral functions at zero temperature the use of several distinct windowing functions has been reported, including Parzen windows [39] and Gaussian windows [28]. Here, we find that a Hann window of the form,
$w(k)=0.5+0.5\cos\left(\frac{\pi k}{S}\right),$ (30)
for $k=0,\ldots,S$ yields the most accurate results when compared to exact diagonalization data in Fig. 4. Hence, we adopt the Hann window as the windowing function in all our subsequent simulations.
Introducing a windowing function reduces the integrated spectral weight. This reduction is commonly remedied by multiplying with a window correction factor [73]. We opt for the so-called energy correction,
$\Gamma=\sqrt{\frac{1}{2S+1}\sum_{k=-S}^{S}w(k)^{2}},$ (31)
conserving the total integrated spectral weight. The estimator for the correct infinite-time Fourier transform is given by,
$\tilde{\mathcal{C}}^{i}(\omega)=\frac{1}{\Gamma}\;\mathcal{C}^{i}(\omega).$ (32)
In the following, we will use the notation $\mathcal{C}^{i}(\omega)$ implicitly for the energy-corrected spectral function $\tilde{\mathcal{C}}^{i}(\omega)$.
III.5 Validation by Exact Diagonalization and DMRG
Figure 5: Dynamical spin structure factor $S^{z}(\bm{q},\omega)$ of the Shastry-Sutherland model at $J/J_{D}=0.63$ on a $16\times 4$ cylinder obtained using the real-time dynamical METTS algorithm at various temperatures in units of the triplon gap $\Delta=0.458J_{D}$. We compare with $S^{z}(\bm{q},\omega)$ at $T=0$ obtained using DMRG for momenta $\bm{q}=(\pi,0)$ (a) and $\bm{q}=(\pi/2,\pi/2)$ (b). Shaded regions depict the standard error of the estimator obtained from $R=100$ METTS states with time evolution up to time $\Omega/J_{D}=50$. MPS time-evolution has been performed using the 2TDVP algorithm with $\chi_{\text{max}}=2000$. We observe that even small features, as well as secondary and tertiary peaks, are well-resolved and converge towards the DMRG result for $T\rightarrow 0$.
To demonstrate the accuracy of our method we start by comparing to data on a $4\times 4$ simulation cluster with periodic boundary conditions, where exact diagonalization (ED) can be straightforwardly performed. In ED, we consider the Lehmann representation of the dynamical spectral function,
$S^{z}(\bm{q},\omega)=\frac{2\pi}{\mathcal{Z}}\sum_{n,m}e^{-\beta E_{n}}|{\braket{m}{S^{z}(\bm{q})}{n}}|^{2}\delta(\omega-(E_{n}-E_{m})),$ (33)
where |$\ket{n}$ and |$\ket{m}$ denotes an eigenbasis with corresponding eigenvalues $E_{n}$ and $E_{m}$ of the Hamiltonian and $\delta$ denotes the Dirac delta function. The eigenvalues and eigenvectors are computed by a full diagonalization of the Hamiltonian matrix. Hence, both the positions of the poles $(\varepsilon_{n}-\varepsilon_{m})$ as well as the weights are computed with machine precision. To analyze and compare the ED spectral function, it is customary to introduce a broadening of the spectral function by replacing the set of Dirac delta peaks with e.g. narrow Gaussian functions of width $\eta$,
$\delta(\omega-p)\quad\longrightarrow\quad\frac{1}{\eta\sqrt{2\pi}}\exp\left[-\frac{(\omega-p)^{2}}{2\eta^{2}}\right].$ (34)
A crucial difference to the proposed METTS method is the fact that the poles and weights are computed exactly and no time evolution is performed. A priori, it is not clear which precision can be expected when only simulating up to a maximal time $\Omega$ and also when only having access to the time-dependent correlator at discrete time steps which are a multiple of $\delta t$. Moreover, it is also not obvious what statistical error we can expect from sampling a finite number of METTS states.
To address these questions, we show results from ED as well as dynamical METTS simulations in Fig. 4. We focus on the parameter $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$ and consider two temperatures $T/J_{D}=0.10$ in (a) and $T/J_{D}=0.20$ (b). We sampled $1000$ METTS and performed a time-evolution up to a maximal time of $\Omega/J_{D}=50$ for every $10$-th METTS. By only considering a subset of all sampled METTS, we verified that the measurement time series has a negligible autocorrelation time. We also employed measurements in the $S^{x}$-basis, to further assure vanishing autocorrelation times, see e.g. [13, 14]. We employed 2TDVP as the time evolution algorithm with a cutoff $\varepsilon=10^{-6}$ without a constraint on the maximal bond dimension $\chi_{\max}$. While a $4\times 4$ simulation cluster is arguably small, we would like to point out that the employed fully periodic boundary conditions are not without challenge for MPS methods, since long-range interactions are present. Hence, this benchmark also serves as a good first test of the quality of time-evolution algorithms employed to compute the METTS states as well as the real-time evolution. To compute the Fourier transform $S(\bm{q},\omega)$ to frequency space, we compared different windowing functions and found the Hann window Eq. 30 to give the most accurate results, cf. Sec. III.4. For ED, we compare three different values of the broadening $\eta=0.02,0.05,0.10$.
We report a close agreement between the ED and METTS simulations. While both the Gaussian broadening $\eta$ for the ED data as well as the finite time horizon $\Omega$ have a broadening effect on the exact spectrum Eq. 33, we observe that even small intricate structures of the spectrum (likely stemming from the small system size) are accurately captured by the METTS algorithm. In particular, the positions of the numerous side peaks are accurately reproduced, as well as the overall spectral amplitude. Choosing $\Omega/J_{D}=50$ and applying the windowing procedure as explained in Sec. III.4 yields a broadening of the spectral function comparable to a Gaussian broadening $\eta\approx 0.05$ in ED. Hence, we conclude that having a finite time horizon $\Omega$ as well as only a discrete set of measurements with time steps $\delta_{t}$ only leads to a small broadening of the spectrum. Moreover, we report that averaging over only $R=100$ METTS yields a satisfyingly small statistical error indicated by the red-shaded region.
We now compare the results from simulations at nonzero temperature with ground state spectral functions from DMRG. This demonstrates the accuracy of the dynamical METTS algorithm in the limit of low temperatures. Results for spectral functions at different temperatures, including $T=0$ from DMRG, are shown in Fig. 5. We focus on the $16\times 4$ cylinder at $J/J_{D}=0.63$ and two wave vectors $\bm{q}=(\pi,0)$ and $\bm{q}=(\pi/2,\pi/2)$. Our maximal time horizon for the real-time evolution has been chosen as $\Omega/J_{D}=50$. Moreover, we collected $R=100$ independent METTS samples for each temperature and observed that this allows sharp error estimates with a large signal-to-noise ratio, as apparent by the small error bars in Fig. 5. The time evolution has been performed using 2TDVP with maximal bond dimensions $\chi_{\max}=1000,1500,2000$. Comparing different bond dimensions for different seeds leads us to conclude that for this particular system, the sample real-time correlation functions $C^{i}(t)$ are already well-converged at $\chi_{\max}=1000$. The spectral function from DMRG has similarly been obtained by a single time-evolution up to time $\Omega/J_{D}=50$. We observe a structured spectrum, where the dominant peak corresponds to single triplon excitations, but also secondary and tertiary peaks corresponding to multi-triplon excitations of the Shastry-Sutherland model. The good convergence towards the DMRG results for $T\rightarrow 0$, where even small features such as secondary and tertiary peaks are well-resolved, demonstrates the accuracy of the approach at low temperatures.
Figure 6: Von Neumann entanglement entropy $S_{\text{vN}}$ of time evolved METTS states $|v_{i}(t)\rangle$ from Eq. 17 on the $16\times 4$ lattice for a bipartition in the center of the lattice. We choose $J/J_{D}=0.63$ and $\bm{q}=(\pi,0)$. Increasing the temperature from $T/J_{D}=0.1$ (a) to $T/J_{D}=0.2$ (b) increases the average entanglement entropy, as shown by the red line with shaded region estimating the statistical error from $R=100$ samples.
A key quantity determining the accuracy of matrix product states is the von Neumann entanglement entropy,
$\mathcal{S}_{\text{vN}}(\rho,A)=-\text{Tr}[\rho_{A}\log(\rho_{A})],$ (35)
where the reduced density matrix with respect to a subsystem $A$ is given by $\rho_{A}=\text{Tr}_{B}[\rho]$ ($B$ denotes the complement of $A$) and $\rho$ denotes an arbitrary density matrix. For pure states |$\ket{\psi}$, the density matrix is given by
$\rho(\ket{\psi})=\ket{\psi}\bra{\psi}.$ | . (36)
For any METTS state |$\ket{\psi_{i}}$, the entanglement entropy in the limit $T\rightarrow\infty$ is vanishing, as then the METTS state is a product state. In the limit $T\rightarrow 0$, the METTS states approach the ground state and its entanglement entropy is naturally bounded. We analyze the behavior of the time-evolved METTS states |$\ket{v_{i}(t)}$ as in Eq. 17 in the dynamical METTS algorithm. Results for the $16\times 4$ cylinder with $J/J_{D}=0.63$ are shown in Fig. 6. The data shown has been obtained using 2TDVP with a maximal bond dimension $\chi_{\max}=2000$. We observe a significant increase in entanglement entropy for a bipartition in the center of the system over the simulated time horizon. Increasing temperature also leads to an increase in entanglement entropy of the time evolved states |$\ket{v_{i}(t)}$ even though the original METTS states |$\ket{\psi_{i}}$ have lower entanglement at a higher temperature. The behavior of the second kind of auxiliary states |$\ket{w_{i}(t)}$ is similar.
Even though the growth of entanglement is sizable for this system, the time-dependent correlation functions $C^{i}(t)$ can still be converged with $\chi_{\max}=1000,1500,2000$ on the width $W=4$ cylinder. This changes when considering cylinders of width $W=6$. There, we found that correlation functions even at a bond dimension $\chi_{\max}=2000$ can be not converged for several samples, especially when increasing temperature. Due to this reason, we introduce the complex-time correlation function as an approach to limit the growth of entanglement in the next section.
IV Complex-time correlation functions
Highly entangled quantum states typically require a large bond dimension to be efficiently compressed using a matrix-product state, which limits the system sizes that can reliably be studied. In the present case, we found the width $W=4$ cylinders to be amenable to the accurate real-time evolution algorithm, whereas $W=6$ is significantly more challenging and ultimately did not lead to converged results given our current computational resources. The main problem hindering such simulations is the entanglement growth with simulation time if performing real-time evolution, as shown in Fig. 6. In the following sections, we discuss how this problem can be ameliorated, by performing complex-time evolution.
IV.1 Two variants of complex-time correlation functions
Real-time evolution of MPS is computationally expensive due to unbounded entanglement growth for generic wave functions. Imaginary-time correlation functions, on the other hand, can be efficiently computed since for imaginary-time $\tau\rightarrow\infty$ every wavefunction $|\phi\rangle$ will converge to the ground state $|\phi\rangle=|\psi_{0}\rangle$ whenever both wave functions have non-vanishing overlap $|\langle\phi|\psi_{0}\rangle|\neq 0$, i.e.
$\lim\limits_{\tau\rightarrow\infty}e^{-\tau H}|\phi\rangle\simeq|\psi_{0}\rangle,$ (37)
where $\simeq$ denotes equivalence up to a phase of the normalized quantum states. In the following, we will parameterize a line through the origin in the complex plane by the angle $\theta$,
$z=t-i\tau=|z|e^{-i\theta}.$ (38)
The negative sign when defining $\tau$ in Eq. 38 is owed to the fact that we aim to perform imaginary time evolutions only with negative exponents and this convention simplifies further notation significantly. For real-time evolution, we write a time-dependent correlation function in terms of the time-evolution operator $U(t)$,
$\mathcal{C}_{AB}(t)=\braket{A(t)B},\text{ where }A(t)=U(t)^{\dagger}AU(t),$ (39)
where $U(t)=\exp(-iHt)$ denotes the unitary time-evolution operator of the time-independent Hamiltonian $H$. $U$ being unitary rests both on the fact that $H$ is hermitian, but also on the fact that $t$ is real. Hence, the complex time-evolution operator of the form,
$\mathcal{U}(z)=e^{-iHz},$ (40)
where $z\in\mathbf{C}$ now denotes a complex time, is generically not a unitary operator for $\text{Im}(z)\neq 0$,
$\mathcal{U}(z)^{-1}\neq\mathcal{U}(z)^{\dagger}.$ (41)
There are two conceivable alternatives when defining complex time correlation functions,
$\mathcal{A}_{AB}(z)$ $=\braket{\mathcal{U}(z)^{-1}A\mathcal{U}(z)B}_{\beta}=\braket{e^{iHz}Ae^{-iHz}B}_{\beta},$ (42)
$\mathcal{H}_{AB}(z)$ $=\braket{\mathcal{U}(z)^{\dagger}A\mathcal{U}(z)B}_{\beta}=\braket{e^{iH\overline{z}}Ae^{-iHz}B}_{\beta}.$ (43)
Here, we denote by $\overline{z}=t+i\tau$ the complex conjugate of $z$. $\mathcal{A}(z)$ is an analytic function of $z$ and therefore we will refer to this quantity as the analytic correlation function. Moreover, for purely imaginary times this correlation function is given by,
$\mathcal{A}_{AB}(-i\tau)=\braket{e^{H\tau}Ae^{-H\tau}B}_{\beta},$ (44)
which is the well-known imaginary-time correlation function used in the Matsubara formalism.
One notices that evaluating $\langle\psi_{i}|e^{H\tau}$ from the left side of Eq. (44) necessarily requires a time evolution process with positive time $\tau$. Generically, computing $e^{+H\tau}|\psi\rangle$ as might occur when evaluating $\mathcal{A}_{AB}(z)$ will ultimately lead to numerical instability. Clearly, this instability issue exists not only for imaginary time evolution but also for generic complex time. As we will demonstrate in Sec. IV.2, this issue in the evaluation of $\mathcal{A}_{AB}(z)$ can be circumvented within a METTS algorithm in an efficient way.
The hermitian correlator $\mathcal{H}_{AB}(z)$ as in Eq. 43, on the other hand, has only recently been investigated in the context of MPS and impurity models [53] and has several intriguing properties which are discussed Sec. IV.4. We choose the name hermitian, since whenever $A$ is a hermitian operator, i.e., $A^{\dagger}=A$, the time-evolved operator $\mathcal{U}(z)^{\dagger}A\mathcal{U}(z)$ is also hermitian. There are computational advantages in evaluating this time-dependent correlation function. For convenience, we again introduce the shorthand notations,
$\mathcal{A}(z)=\mathcal{A}_{AB}(z),$ (45)
$\mathcal{H}(z)=\mathcal{H}_{AB}(z).$ (46)
Figure 7: Sketch of the complex contour on which the analytical correlation function $\mathcal{A}(z)$ is evaluated. The maximal imaginary time is chosen to be $\Omega_{\tau}=\beta/2$ in order to allow for an efficient and stable numerical algorithm.
IV.2 Analytic complex-time correlation functions
We propose an efficient method for evaluating the analytic complex time correlation function $\mathcal{A}(z)$. Its definition as in Eq. 42 is expressed as,
$\mathcal{A}(z)$ $=\braket{e^{iHz}Ae^{-iHz}B}_{\beta}$ (47)
$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\langle\psi_{i}|e^{izH}Ae^{-izH}B|\psi_{i}\rangle$ (48)
into the basis of METTS |$\ket{\psi_{i}}$. We choose to evaluate the analytical correlator on the contour $z=|z|e^{-i\theta}$ for,
$|z|\in[0,\Omega],\quad\text{Re}(z)\geq 0,$ (49)
where $\Omega$ denotes a final time horizon. We define,
$\Omega_{t}=\Omega\cos\theta,\quad\Omega_{\tau}=\Omega\sin\theta,\quad\Omega_{z}=\Omega_{t}-i\Omega_{\tau}.$ (50)
Thus, the real and imaginary time components take values,
$t\in[0,\Omega_{t}]\subseteq\mathbb{R},\quad\tau\in[0,\Omega_{\tau}]\subseteq\mathbb{R}.$ (51)
As shown further below, when choosing
$\Omega_{\tau}=\frac{\beta}{2}$ (52)
the analytic correlator $\mathcal{A}(z)$ can be evaluated without ever having to perform an imaginary time evolution with positive time $e^{+\tau H}\ket{\psi}$. For actual computations, we discretize the complex time $z$ into $S$ steps with step size $\delta_{z}=\Omega_{z}/S$, i.e.
[…]
where,
$\delta_{z}=\delta_{t}-i\delta_{\tau}=|\delta_{z}|e^{-i\theta}.$ (54)
The complex contour and the quantities $\Omega_{z}$, $\Omega_{t}$, $\Omega_{\tau}$ for which we evaluate $\mathcal{A}(z)$ are shown in Fig. 7.
A naive approach to evaluating $\mathcal{A}(z)$ would be to compute intermediate states,
$|v_{i}(t)\rangle=e^{-iHz}B|\psi_{i}\rangle=e^{-iHt}e^{-\tau H}B|\psi_{i}\rangle$ (55)
$|w_{i}(t)\rangle=e^{-iH\overline{z}}|\psi_{i}\rangle=e^{-iHt}e^{+\tau H}|\psi_{i}\rangle,$ (56)
with $z=t-i\tau$ similar to the real-time correlation algorithm Alg. 1. However, this approach is plagued by severe numerical instability when computing the state $|w_{i}(t)\rangle$. Computing a positive exponential $e^{+\tau H}|\psi_{i}\rangle$ is numerically highly unstable and prohibits most practical calculations.
When restricting our time horizon to $\Omega_{\tau}=\beta/2$ we can circumvent ever having to compute positive real exponential in the context of the METTS algorithm. The key idea is that the positive imaginary time evolution can be absorbed by the negative imaginary time evolution performed when calculating an individual METTS |$\ket{\psi_{i}}$. For this, we rewrite Eq. 47 as,
Algorithm 2 Dynamical METTS for analytic correlator $\mathcal{A}(z)$
1:Choose a random initial product state $|\sigma_{1}\rangle=|\sigma_{1}^{1}\rangle\cdot\ldots|\sigma_{1}^{N}\rangle$
2:for $i=1,\ldots,R$ do
3:     Compute $|\psi_{i}\rangle=\frac{1}{\sqrt{p_{i}}}e^{-\beta H/2}|\sigma_{i}\rangle$ and $p_{i}=\braket{\sigma_{i}}{e^{-\beta H}}{\sigma_{i}}$
4:     Set |$\ket{\phi_{i}^{0}}=\frac{1}{\sqrt{p_{i}}}\ket{\sigma_{i}}$
5:     for k  =  1, …, S do
6:         Compute |$\ket{\phi_{i}^{k}}=e^{i\delta_{t}H}e^{-\delta_{\tau}H}\ket{\phi_{i}^{k-1}}$
7:         Store |$\ket{\phi_{i}^{k}}$ to disk for later use
8:     end for
9:     for k = 0, …, S do
10:         if $k=0$ then
11:              Compute $|\chi_{i}^{0}\rangle=B|\phi_{i}^{S}\rangle$
12:         else
13:              Compute |$\ket{\chi_{i}^{k}}=e^{-i\delta_{t}H}e^{-\delta_{\tau}H}\ket{\chi_{i}^{k-1}}$
14:         end if
15:         Read |$\ket{\phi_{i}^{S-k}}$ from disk
16:         Evaluate $\mathcal{A}^{i}(k\delta_{z})=\bra{\phi^{S-k}_{i}}A\ket{\chi_{i}^{k}}$
17:     end for
18:     Pick next product state $|\sigma_{i+1}\rangle$ with probability $|\langle\psi_{i}|\sigma_{i+1}\rangle|^{2}$
19:end for
$\mathcal{A}$ $(z)=\braket{e^{izH}Ae^{-izH}B}_{\beta}=$ (57)
$=\braket{e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}}_{\beta}$ (58)
$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle$ (59)
$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\;\mathcal{A}^{i}(z),$ (60)
where we defined the sample analytic correlation function as,
$\mathcal{A}^{i}(z)\equiv\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle.$ (61)
Introducing the factor $e^{\pm i\Omega_{t}H}$ as in Eq. 61 allows to rewrite $\mathcal{A}^{i}(z)$ in a simple form,
$\mathcal{A}^{i}(z)$ $=\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle$ (62)
$=\frac{1}{p_{i}}\langle\sigma_{i}|e^{-i(\Omega_{z}-z)H}Ae^{-izH}Be^{i\overline{\Omega}_{z}H}|\sigma_{i}\rangle.$ (63)
For $k=0,\ldots,S$ we introduce the auxiliary computational states |$\ket{\phi_{i}^{k}}$ and |$\ket{\chi_{i}^{k}}$ defined as,
|$\ket{\phi_{i}^{k}}$ $\equiv\frac{1}{\sqrt{p_{i}}}e^{ik\overline{\delta}_{z}H}\ket{\sigma_{i}}$ $=\frac{1}{\sqrt{p_{i}}}e^{ik{\delta_{t}}H}e^{-k{\delta_{\tau}}H}\ket{\sigma_{i}},$ (64)
|$\ket{\chi_{i}^{k}}$ $\equiv e^{-ik\delta_{z}H}B\ket{\phi_{i}^{S}}$ $=e^{-ik\delta_{t}H}e^{-k\delta_{\tau}H}B\ket{\phi_{i}^{S}}.$ (65)
Importantly, we see that these states can all be computed by performing a negative imaginary time evolution, i.e., computing $e^{-\tau H}\ket{\psi}$ for some state |$\ket{\psi}$. This insight is the main reason why $\mathcal{A}^{i}(z)$ can be computed efficiently. Notice that,
$\bra{\phi_{i}^{S-k}}$ | $=\frac{1}{\sqrt{p_{i}}}\bra{\sigma_{i}}e^{-i(\Omega_{z}-k\delta_{z})H},$ (66)
|$\ket{\phi_{i}^{S}}$ $=\frac{1}{\sqrt{p_{i}}}e^{i\overline{\Omega}_{z}H}|\sigma_{i}\rangle.$ (67)
Using a discretized $z=k\delta_{z}$ we can rewrite Eq. 63 as,
$\mathcal{A}^{i}(k\delta_{z})=\bra{\phi^{S-k}_{i}}A\ket{\chi_{i}^{k}}.$ (68)
To evaluate these expectation values for all discretized values of $z=k\delta_{z}$ ($k=0,\ldots,S$), we compute the states |$\ket{\phi_{i}^{k}}$ for $k=0,\ldots,S$ successively by complex time evolution and store all intermediate states to disk for later use. Then we compute the state |$\ket{\chi^{0}_{i}}=B\ket{\psi_{i}^{S}}$ from which we successively compute |$\ket{\chi^{k}_{i}}$ for $k=1,\ldots,S$ and thereby evaluate the desired correlation functions $\mathcal{A}^{i}(z)=\bra{\phi^{S-k}_{i}}A\ket{\chi_{i}^{k}}$. $\mathcal{A}(z)$ is then obtained by averaging over $\mathcal{A}^{i}(z)$. We summarize the described algorithm in Alg. 2.
Alg. 2 is in several aspects similar to Alg. 1. Both algorithms require computing the METTS |$\ket{\psi_{i}}$ and then time-evolves two further states to compute time-dependent correlators. Moreover, the basic thermal METTS Markov chain is constructed in both cases. A difference in Alg. 2 is the necessity of storing the auxiliary states |$\ket{\phi_{i}^{k}}$. On present-day computers, this typically neither imposes severe time nor memory constraints. However, in the limit of small angle $\theta$ the memory requirement can be considerable due to the large number of discrete time steps and requires corresponding storage capacities.
The key quantity determining the efficiency of the algorithm is the bipartite von-Neumann entanglement entropy $S_{\text{vN}}$ as in Eq. 35 of the MPS |$\ket{\phi_{i}^{k}}$ and |$\ket{\chi_{i}^{k}}$, shown in Figs. 8 and 9, respectively. The entanglement is considered for a bipartition at the center of the system. We evaluated the analytic correlator for the dynamical spin structure factor $S^{z}(\bm{q},\omega)$ as in Eq. 2 for the $16\times 4$ Shastry-Sutherland model at $J/J_{D}=0.63$ at temperatures $T/J_{D}=0.1$ and $T/J_{D}=0.2$. The auxiliary state |$\ket{\phi_{i}^{0}}=\ket{\sigma_{i}}$ is a classical product state such that the entropy in Fig. 8 starts from zero. Slower entanglement growth is observed for larger values of $\theta$ as expected. For $\theta=\pi/2$ we are performing fully imaginary time evolution, which yields the slowest entanglement growth. Meanwhile, fluctuations of $S_{\text{vN}}$ among the sampled states during METTS dynamics are rather small, as indicated by the shaded regions around the mean. The absolute values of $S_{\text{vN}}$ are comparable for both |$\ket{\phi_{i}^{k}}$ and |$\ket{\chi_{i}^{k}}$ and grow with increasing temperature. We observe a significant reduction in entanglement entropy when comparing the values to the entanglement of the states $|v_{i}(t)\rangle$ Eq. 17 as shown in Fig. 6. While we observed maximal values of $S_{\text{vN}}\approx 5.3$ for $T/J_{D}=0.2$ for $|v_{i}(t)\rangle$ in Fig. 6, we only find maximal values of $S_{\text{vN}}\approx 2.3$ for $\theta=\pi/8$ in Figs. 8 and 9. Naturally, the chosen time horizon $\Omega$ is significantly shorter when evaluating the analytical correlator. However, this demonstrates that $\mathcal{A}(z)$ can be evaluated at significantly reduced cost at non-zero values of $\theta$. We want to point out that setting a value of $\theta$ determines the value of $\Omega$ due to the constraint $\Omega_{\tau}=\beta/2$ we impose for the analytical correlator. Hence, the maximal complex times $|z|$ differ for different values of $\theta$.
Figure 8: Average bipartite von Neumann entanglement entropy $S_{\text{vN}}$ of the auxiliary state |$\ket{\phi_{i}^{k}}$ Eq. 64 for evaluating the analytical correlation function $\mathcal{A}(z)$. We consider the $16\times 4$ Shastry-Sutherland lattice for different values of the complex angle $\theta$ and $J/J_{D}=0.63$ and compare results for temperatures (a) $T/J_{D}=0.1$ and (b) $T/J_{D}=0.2$. The average is computed from $R=100$ METTS samples and the shaded region shows the estimated standard deviation of the mean.
Figure 9: Same as Fig. 8, but for the auxiliary state |$\ket{\chi_{i}^{k}}$ Eq. 65.
IV.3 Stochastic analytic continuation
Figure 10: Dynamical structure factors ${S}(\bm{q},\omega)$ from stochastic analytic continuation based on analytic correlation functions $S(\bm{q},z)$ at several complex angles. Here the spectrum in obtained from sampling at $\alpha=\alpha_{c}$. Data is shown for the $16\times 4$ lattice with $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$. We compare between temperatures $T/J_{D}=0.1$ (a) and $T/J_{D}=0.2$ (b).
Figure 11: Same as Fig. 10, where the spectrum is obtained via weighted averaging the collected histogram from $\log\alpha=0$ to $\log\alpha=-6$.
To retrieve the spectral function from $\mathcal{A}(z)$ we perform analytic continuation, for which various approaches have been proposed and tested in the context of quantum Monte Carlo simulations [74, 75, 76, 77]. Although mostly used for performing continuation from imaginary time correlation functions, these approaches are often also applicable for the analytic correlator $\mathcal{A}(z)$. Generally, inverse Laplacian approaches from imaginary time correlation functions lead to a biased result where the details of the efficiency are case-dependent. In most cases, dominant peaks capturing well-defined quasiparticles can be resolved in a satisfactory way. However, for a spectrum with contributions from both a low energy peak and a higher energy continuum, numerical continuation is hard to quantitatively resolve the information at high frequency. On the other hand, we propose the numerical analytic continuation to be more efficient via correlation functions defined on the complex time case, since real-time contributions of the correlator, even for short periods of time, carry information about the high-frequency spectrum.
We now document our approach to stochastic analytic continuation (SAC) based on correlation functions in the complex time plane. In the following, we choose the observables to be $A^{\dagger}=B\equiv\mathcal{O}$. The aim is to obtain the finite temperature spectral function,
[…]
where $|m\rangle$ ($|n\rangle$) is the $m$-th ($n$-th) eigenvector of the Hamiltonian with energy $E_{m}$ ($E_{n}$) cf. Eq. 33.
The correlation function is related to the spectral function via the following Laplacian transformation,
[…]
where $S(-\omega)=e^{-\beta\omega}S(\omega)$ and we defined $A(\omega)\equiv S(\omega)(1+e^{-\beta\omega})/2$ and the integration kernel as,
$K(z,\omega)\equiv(e^{-z\omega}+e^{-(\beta-z)\omega})(1+e^{-\beta\omega})^{-1}\pi^{-1},$ (71)
which is by definition also a complex function. We parameterize $A(\omega)$ as a sum of delta functions,
$A(\omega)=\sum_{i}^{N}A_{i}\delta(\omega-\omega_{i}).$ (72)
During the stochastic sampling process, the testing Green’s function is,
[…]
We run stochastic sampling by shifting the positions and amplitudes of the delta functions in Eq. 72, based on the following probability distribution,
$P(A)\propto e^{-\frac{\chi^{2}(A)}{\alpha}},$ (74)
where $\alpha$ denotes the fictitious inverse temperate, and $\chi^{2}(A)$ is the $\chi^{2}$ that measures the deviation of correlation function based on any given $A(\omega)$ during stochastic sampling, with the physical one that is observed from METTS calculation. In the current case, fluctuations from both real and imaginary parts of the correlation function are included in the definition of $\chi^{2}$,
$\chi^{2}\equiv\sum_{k=1}^{S}\sum_{k^{\prime}=1}^{S}$ $\left({C^{R}_{kk^{\prime}}}^{-1}\text{Re}(\mathcal{A}^{S}(z_{k})-\overline{\mathcal{A}}(z_{k^{\prime}}))^{2}\right.$ (75)
$+$ $\left.{C^{I}_{kk^{\prime}}}^{-1}\text{Im}(\mathcal{A}^{S}(z_{k})-\overline{\mathcal{A}}(z_{k^{\prime}}))^{2}\right),$
where $\overline{\mathcal{A}}(z_{k})$ denotes the mean value of the analytic correlator at the $k$-th time step $z_{k}=k\delta_{z}$. The covariance matrices corresponding to the real and imaginary parts of the correlator are defined respectively,
$C^{R}_{kk^{\prime}}\equiv$ $\frac{1}{R(R-1)}$ (76)
$\sum_{i=1}^{R}\text{Re}(\mathcal{A}^{i}(z_{k})-\overline{\mathcal{A}}(z_{k}))\text{Re}(\mathcal{A}^{i}(z_{k^{\prime}})-\overline{\mathcal{A}}(z_{k^{\prime}}))$
$C^{I}_{kk^{\prime}}\equiv$ $\frac{1}{R(R-1)}$
$\sum_{i=1}^{R}\text{Im}(\mathcal{A}^{i}(z_{k})-\overline{\mathcal{A}}(z_{k}))\text{Im}(\mathcal{A}^{i}(z_{k^{\prime}})-\overline{\mathcal{A}}(z_{k^{\prime}})),$
where $\mathcal{A}^{i}(z_{k})$ is the correlation function measured from the $i$-th METTS configuration.
So far, the major difference between the current setup and the standard analytic continuation approach based on imaginary time correlation functions is the complexity of time $z$, correlation function $G(z)$, and the integration kernel $K(z,\omega)$. The determination of the optimal fictitious sampling temperature $\alpha$ is a crucial issue in SAC methods. It relies on the dependence of the average $\chi^{2}$ on temperature $\alpha$, which is case-dependent. The crossover regime where $\chi^{2}$ gradually saturates to the minimum value of chi-squared $\chi^{2}_{\text{min}}$ is usually argued to be the optimal choice. On the other hand, in the case that the time domain is far from the imaginary axis, we observe a mild decay of $\chi^{2}$ as $\alpha$ is decreased from $1$. Here we test two approaches to evaluating the spectrum. The first method is to perform an annealing process sequentially from $\alpha=1$ to low temperature and pick up the optimal sampling temperature $\alpha=\alpha_{c}$ as the one where the corresponding $\chi^{2}$ satisfies,
$\chi^{2}=\chi^{2}_{\text{min}}+\sqrt{2S}$ (77)
and we collect the averaged spectrum $\alpha_{c}$. The second approach is to average the sampled spectrum at a temperature list $\ln(1/\alpha)=0,0.1,0.2,...6.0$ via annealing. Note that this is a weighted average and the weight is proportional to the slope of $\chi^{2}$-$\alpha$ curve. Here, the formal approach is a simple generalization of the one used in Ref. [77] to the complex plane, whereas the latter method, which was also frequently used in imaginary time cases, follows Ref. [76].
We show results of analytic continuation at $T/J_{D}=0.1$ and $0.2$ for several complex angles in Figs. 10 and 11. First, the spectrum obtained via the sampling at $\alpha_{c}$ does not differ from the one via weighted averaging the data from high to low temperature. Clearly, the position of the lowest triplon mode at $T/J_{D}=0.1$ is constant with both perturbation theory and the one from hermitian correlation function (at $\theta\longrightarrow 0$ limit) as shown in Fig. 12. Meanwhile, the second triplon peak at a higher frequency, which is not clearly visible for the case of $\theta=\pi/2$, can be resolved better from complex angle calculations. On the other hand, we did not observe a significant $\theta$ dependence for the thermal broadened spectrum at $T/J_{D}=0.2$.
IV.4 Hermitian complex-time correlation functions
Figure 12: Hermitian dynamical structure factors $\tilde{S}^{z}_{\theta}(\bm{q},\omega)$ from complex-time correlation functions $\tilde{S}^{z}_{\theta}(\bm{q},t)$ at different complex angles $\theta$. Data is shown for the $16\times 4$ lattice with $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$. We compare between temperatures $T/J_{D}=0.1$ (a) and $T/J_{D}=0.2$ (b). We observe that several key features of the exact spectral function $S(\bm{q},\omega)$ ($\theta=0$) are already captured well by sizable values of $\theta$. Larger values of $\theta$ significantly reduce the computational cost. Time evolution has been performed up to time $\Omega/J_{D}=50$ for $R=100$ samples.
We now focus our attention on the second possible generalization of correlation functions to complex time, the hermitian correlation function as defined in Eq. 43, which expanded in the METTS basis yields,
$\mathcal{H}_{AB}(z)$ $=\braket{e^{iH\overline{z}}Ae^{-iHz}B}$ (78)
$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\langle\psi_{i}|e^{iH\overline{z}}Ae^{-iHz}B|\psi_{i}\rangle$ (79)
$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\mathcal{H}_{AB}^{i}(z),$ (80)
where we define the sample hermitian correlation function as,
$\mathcal{H}^{i}_{AB}(z)\equiv\langle\psi_{i}|e^{iH\overline{z}}Ae^{-iHz}B|\psi_{i}\rangle.$ (81)
To evaluate this correlation function we can proceed analogously to the conventional time-dependent correlation function algorithm in Alg. 1, and define the states,
$|v_{i}(z)\rangle=e^{-iHz}B|\psi_{i}\rangle=e^{-H\tau}e^{-iHt}B|\psi_{i}\rangle,$ (82)
$|w_{i}(z)\rangle=e^{iH\overline{z}}|\psi_{i}\rangle=e^{-H\tau}e^{-iHt}|\psi_{i}\rangle.$ (83)
Using these states, we evaluate the sample hermitian correlation functions as,
$\mathcal{H}^{i}_{AB}(z)=\langle w_{i}(z)|A|v_{i}(z)\rangle,$ (84)
using Alg. 1 adjusted for the complex time evolution along the contour $z=|z|e^{-i\theta}$ shown in Fig. 7. A key difference to the dynamical METTS algorithm Alg. 1 is the fact that the time evolution operators $e^{-iHz}$ and $e^{iH\overline{z}}$ are not norm preserving. However, we could reinforce a normalization by setting,
$n_{v}^{2}$ $\equiv\braket{v_{i}(z)}{v_{i}(z)}=\bra{\psi_{i}}B^{\dagger}e^{-2\tau H}B\ket{\psi_{i}},$ (85)
$n_{w}^{2}$ $\equiv\braket{w_{i}(z)}{w_{i}(z)}=\bra{\psi_{i}}e^{-2\tau H}\ket{\psi_{i}},$ (86)
$c^{2}_{B}$ $\equiv\braket{\psi_{i}}{B^{\dagger}B}{\psi_{i}}$ (87)
and introducing the modified states,
$|\tilde{v}_{i}(z)\rangle=c_{B}\;|\tilde{v}_{i}(z)\rangle/n_{v}=c_{B}\;e^{-iHt}e^{-H\tau}B|\psi_{i}\rangle/n_{v},$ (88)
$|\tilde{w}_{i}(z)\rangle=|\tilde{w}_{i}(z)\rangle/n_{w}=e^{-H\tau}e^{-iHt}e^{-H\tau}|\psi_{i}\rangle/n_{w}.$ (89)
Finally, we consider the modified sample hermitian correlation functions,
~$\tilde{\mathcal{H}}^{i}_{AB}(z)=\langle\tilde{w}_{i}(z)|A|\tilde{v}_{i}(z)\rangle,$ (90)
giving rise to a modified hermitian correlation function
~$\tilde{\mathcal{H}}_{AB}(z)=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\tilde{\mathcal{H}}^{i}_{AB}(z).$ (91)
Our strategy is to directly view ~$\tilde{\mathcal{H}}_{AB}(z)$ as an approximation of $\mathcal{C}_{AB}(t)$,
$\mathcal{C}_{AB}(t)\approx\tilde{\mathcal{H}}_{AB}(t-it\tan\theta).$ (92)
We find that ~$\tilde{\mathcal{H}}_{AB}(z)$ is a significantly better approximation to $\mathcal{C}_{AB}(t)$ than the bare hermitian correlator $\mathcal{H}_{AB}(z)$, which motivates the modified definition.
The negative signs in the real exponentials $e^{-H\tau}$ have significantly favorable consequences on the stability of computing $e^{-H\tau}|\psi\rangle$, where $|\psi\rangle$ denotes an arbitrary wave function. Moreover, evaluating imaginary-time evolutions $e^{-H\tau}|\psi\rangle$ with a negative prefactor is well-suited for MPS methods. Whenever the overlap with the groundstate |$\ket{\psi_{0}}$ of $H$ is nonzero, i.e., $\braket{\psi_{0}}{\psi}\neq 0$, we have
$\lim_{\tau\rightarrow\infty}e^{-H\tau}|\psi\rangle\simeq\ket{\psi_{0}}.$ (93)
Since ground states of local Hamiltonians are subject to bounds on their entanglement, we also expect states of the form $e^{-H\tau}|\psi\rangle$ to have manageable entanglement, whenever $|\psi\rangle$ is weakly entangled.
We show hermitian spectral functions $S^{z}_{\theta}(\bm{q},\omega)$ for the $16\times 4$ Shastry-Sutherland lattice at $J/J_{D}=0.63$ and temperatures $T/J_{D}=0.1,0.2$ in Fig. 12. Remarkably, for $T/J_{D}=0.1$ the position of the dominant peak at $\omega/J_{D}\approx 0.46$ is always obtained consistently even for sizable values of the complex angle $\theta$. Meanwhile, the height and width of the dominant peak are well captured as long as $\theta\leq\pi/16$. Moreover, even features like secondary peaks at higher temperatures are captured correctly for not too large values of the complex angle $\theta$. Similarly, the overall shape of the exact spectral function with $\theta=0$ at $T/J_{D}=0.2$ in Fig. 12(b) is generally well captured at larger values of $\theta$. However, the peak around the triplon gap for $\omega/J_{D}\approx 0.46$ is overestimated at finite angles of $\theta$.
Figure 13: Average bipartite von Neumann entanglement entropy $S_{\text{vN}}$ of the states |$\ket{\tilde{v}_{i}(z)}$ as in Eq. 88 for computing the hermitian correlator Eq. 43. Data on the $16\times 6$ Shastry-Sutherland lattice for different values of the complex angle $\theta$ at $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$. We compare results for temperatures $T/J_{D}=0.1$ (a) and $T/J_{D}=0.2$ (b). For larger values of $\theta$ we observe convergence of $S_{\text{vN}}$ towards the ground state entanglement entropy.
Our main motivation for employing the hermitian correlation function is to limit the entanglement when time-evolving an MPS state. Thus, we show the average growth of the bipartite von Neumann entropy $S_{\text{vN}}$ of the states |$\ket{\tilde{v}_{i}(z)}$ for different values of $\theta$ on a $16\times 6$ cylinder in Fig. 13. We chose momenta $\bm{q}=(\pi/2,\pi/2)$, $J/J_{D}=0.63$ and temperatures $T/J_{D}=0.1,0.2$. As expected, for $\theta=0$ the entanglement is monotonically growing as a function of time. While the initial entanglement of the METTS state at $T/J_{D}=0.1$ in (a) is larger than at $T/J_{D}=0.2$ in (b), the entanglement grows more quickly when going to higher temperatures. As soon as $\theta>0$ we observe an initial growth of the entanglement, followed by a decay caused by the imaginary-time evolution component when applying $e^{-\tau H}$. We see at nonzero values of $\theta$ that entanglement is quickly suppressed after attaining an intermediate maximum which grows when decreasing $\theta$. We thus conclude that a nonzero complex angle $\theta$ solves the issue of unbounded entanglement growth at the cost of introducing a controlled error when evaluating the spectral functions.
V Discussion and Perspectives
[…]
The real-time evolution of METTS states performed is the computational bottleneck due to the generically unbounded growth of entanglement. To alleviate the difficulties arising from large bond dimensions, we suggest that complex-time correlators can be used to lower the computational cost. The key idea is that a contribution from imaginary-time evolution with a negative exponent will ultimately evolve the METTS state to the ground state, for which entanglement bounds are known. While complex-time evolution had been proposed for QMC [51] or MPS methods at $T=0$ applied to impurity models [52, 53], we here propose the algorithms to employ this approach at non-zero temperatures using METTS with MPS. We considered two different ways of generalizing time-dependent correlation functions to complex time which we denote by the analytic and hermitian correlation functions. The analytic one is a natural generalization of the spectrum in Matsubara formalism to the complex plane, such that conventional numerical analytic continuation approaches can be easily generalized correspondingly. Most importantly, real-time contribution to the time-displaced correlator improves the quality of analytic continuation. Ref. [52] also provided an alternative method for performing the continuation from the complex plane, which could presumably also be applied to the present case.
On the other hand, the definition of the hermitian correlator has the advantage of easy implementation and the imaginary time evolution dissipates the entanglement entropy dramatically. Meanwhile, the main structure of the real frequency spectrum is well captured even in the presence of a systemic deviation due to imaginary contributions. Only recently, this version of a complex time correlator has been studied [53] in the context of impurity models with MPS.
It is worth pointing out, that the idea of introducing an imaginary-time component to reduce entanglement is similar in spirit to the dissipation-assisted operator evolution proposed in Ref. [45]. The imaginary-time component acts as a natural ”dissipator”. One advantage of the present approach is that existing implementations of time-evolution algorithms, such as TDVP, can be readily employed to compute complex-time evolution.
The SAC method based on the complex plane correlation function has still leftover possibilities to be improved in the future: one promising suggestion is to perform continuation via combined information of correlators at several angles along the complex plane. Apart from the stochastic approach, an interesting direction is to combine the Nevanlinna analytic continuation [78] approach with the correlation functions obtained at certain METTS snapshots. On the other hand, a common technique for computing dynamical spectral functions at zero temperatures by using real-time evolution with MPS, is to perform linear prediction in order to extend the time-series of correlation functions [79, 39, 80, 81]. While it is reported that this works remarkably well for simple spectra with few poles, we have not been able to successfully apply this to the correlation functions computed at non-zero temperatures in this work. For generic METTS states we mostly find strong disagreement between the actual time-evolution and the linear prediction. This is likely due to the fact that the correlators $\mathcal{C}_{i}(t)$ from METTS in general have a more complex structure than a superposition of a few sinusoidal waves. Thus, all the data we present is taken from actual time evolution. Recently, however, a more advanced method for performing linear prediction in the context of spectral functions with increased accuracy has been proposed [81]. It would be interesting to find out whether this method would be beneficial in the present context to extend the METTS time series.
In summary, we demonstrated that the combination of METTS with either real or complex-time evolution can be applied to study spectral functions of two-dimensional quantum many-body systems using MPS. While the most accurate approach using real-time correlation functions is feasible on certain problem sizes, the complex-time correlation functions can be evaluated more efficiently. The reduction of computation cost can be tuned using the complex angle $\theta$, from highly accurate but less efficient at $\theta$ close to zero to highly efficient but less accurate at larger values of $\theta$. We, therefore, propose this method to serve as a bridge between the physics quantum lattice models and spectroscopy experiments performed on strongly correlated materials.
Acknowledgements.
Z.W. thanks Anders Sandvik and Hui Shao for enlightening discussions on the stochastic analytic continuation method. A.W. is grateful to Siddharth Parameswaran, Roderich Moessner, Miles Stoudenmire, and Olivier Parcollet for insightful discussion. A.W. acknowledges support by the DFG through the Emmy Noether program (Grant No. 509755282). Z.W. was supported by the FP7/ERC Consolidator Grant QSIMCORR, No. 771891 and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2111 – 390814868.
Appendix A Projection error of TDVP
Figure 14: Imaginary time correlation function using TEBD and the 2-site TDVP approach, for $T/\Delta=0$ (a), $0.1$ (b) and $0.2$ (c). Two dashed lines, including $c_{2}e^{-\Delta\tau}$ ( $\Delta\approx 0.459$ ) and $c_{1}e^{-0.56\tau}$ are guides to the eye. The time evolution approach for measuring correlation function at $T=0$ is based on the ground state via DMRG calculation, where the ones at $T/J_{D}=0.1$ and $0.2$ are from averaging the METTS configurations.
One potential error of the TDVP method exists at the step of projecting the time-dependent Schrödinger equation onto the relevant MPS subspace [67]. This projection error is especially severe when time evolving an MPS with very low bond dimension. It turns out to be a serious issue for the Shastry-Sutherland model when the temperature is much lower than the first triplon gap $\Delta$. In this low energy case, the MPS during the time evolution step does not differ significantly from the ground state, which is known to be a product state of spin singlets living on the off-diagonal bonds. The projection error leads to a biased estimation of the spin gap, even with the 2-site TDVP approach. On the other hand, time-evolving block decimation (TEBD) suffers only from the Trotter error and the MPS truncation error.
As a test, we show a comparison between TDVP and TEBD methods in Fig. 14. At low temperatures, one expects the imaginary time correlation function to decay with an exponential factor based on the estimated spin gap ($\Delta\approx 0.459J_{D}$) from the perturbation theory. However, as can be seen from the plot, $G(\tau)$ obtained from the 2-site TDVP method deviates from this expectation with a much steeper slope, for $T/\Delta=0$ and $0.1$. On the other hand, the long-time correlation from TEBD resolves the expected spin gap. Meanwhile, for temperatures comparable with the spin gap, enough entanglement is built up within the time-evolved MPS and we do not observe a severe projection error in TDVP. This can be seen from the consistent data between the two methods at $T/\Delta=0.2$. For the three temperatures here, we choose the time step in TEBD as $d\tau=0.1$ such that the Trotter error is sufficiently small.
Appendix B Fictitious temperature dependence of stochastic analytic continuation
In the SAC sampling process, one normally expects a characteristic thermal phase transition temperature where the specific heat shows a jump. However, in the SAC process based on our METTS data on the $16\times 4$ Shastry-Sutherland lattice, $\chi^{2}$ decays toward $\chi^{2}_{\text{min}}$ smoothly as a function of the fictitious temperature $\alpha$. Hence we choose to use the weighted average of the spectrum over a wide range of $\alpha$. It is nevertheless important to check the $\alpha$ dependence of the spectrum, and especially to compare between the imaginary and complex time cases.
Figure 15: ${S}(\bm{q},\omega)$ from stochastic analytic continuation at $\alpha=1.0,0.1,0.01$ and $0.0045$, based on ${S}(\bm{q},z)$ at $\theta=\pi/2$ (a) and $\pi/8$ (b), for $T/J_{D}=0.1$. The simulations at different $\alpha$ are performed via independent annealing processes from high temperatures.
We choose the same set of METTS data that is used in Sec. IV.3 as showcased here. As depicted in Fig. 15, the spectral function obtained via the imaginary time $\theta=\pi/2$ correlator depends strongly on the choice of $\alpha$. Taking the spectrum in Fig. 12 as a reference, the weight and position of the second triplon peak are clearly biased for both $\alpha=0.01$ and $0.0045$. On the other hand, the overall shape of the spectrum at $\theta=\pi/8$ is much more stable, although we observe an unphysical splitting of the second peak at small $\alpha$.

### v2

@@
 Recently, however, complex time correlation functions have been revisited in the context of MPS ground state simulations and applied to quantum impurity models [52, 53].
-As we will show, complex-time correlation functions can be evaluated using well-known MPS time evolution algorithms such as the time-dependent variation principle (TDVP) [54, 55].
+As we will show, complex-time correlation functions can be evaluated using well-known MPS time evolution algorithms such as the time-dependent variational principle (TDVP) [54, 55].
 Our aim is to combine the advantages of computing imaginary-time correlations and real-time correlations.
@@
 […]
+$\Delta=J_{D}\left[1-(J/{J_{D}})^{2}-\frac{1}{2}(J/{J_{D}})^{3}-\frac{1}{8}(J/{J_{D}})^{4}\right].$ (1)
+[…]
 […]
@@
 We begin with Sec.
-II introducing the Shastry-Sutherland model we use to benchmark out methods.
+II introducing the Shastry-Sutherland model we use to benchmark our methods.
 In Sec.
@@
 Flipping such a dangling spin would constitute an artificial boundary mode and, therefore, we remove interdimer couplings that couple to these boundary dangling spins.
-Figure 2: Illustration of the dynamical METTS algorithm.
-A sequence of product states $|\sigma_{i}\rangle$ and METTS $|\psi_{i}\rangle$ is sampled according to the thermal METTS algorithm.
-For every METTS the time-dependent correlation function $\mathcal{C}_{AB}^{i}(t)$ as in Eq.
-20 is computed.
-Product states $|\sigma_{i+1}\rangle$ are sampled with transition probability $T_{i\rightarrow i+1}=|\langle\psi_{i}|\sigma_{i+1}\rangle|$.
-From the individual METTS correlation functions $\mathcal{C}_{AB}^{i}(t)$ the spectral function $\mathcal{C}_{AB}(\omega)$ is computed via a Fourier transform and subsequent averaging.
 We are interested in the dynamical spin structure factor given by
@@
 1(b).
+he Shastry-Sutherland model features a flat band of triplon excitations, which stretches throughout the entire Brillouin zone.
+At low temperatures we expect similar physical phenomena to occur regardless of the exact momentum.
+In this manuscript we focus on two momenta, $\bm{q}=(\pi,0)$ and $\bm{q}=(\pi/2,\pi/2)$ to illustrate the physics, but expect similar behavior at different values of $\bm{q}$.
 To reduce the boundary effects of the cylinder we use a modified definition,
@@
 $|\psi_{i}\rangle=\frac{1}{\sqrt{p_{i}}}e^{-\beta H/2}|\sigma_{i}\rangle,$ (11)
-where $p_{i}=\langle\sigma_{i}|e^{-\beta H}|\sigma_{i}\rangle\geq 0$ is is a (non-negative) weight.
+where $p_{i}=\langle\sigma_{i}|e^{-\beta H}|\sigma_{i}\rangle\geq 0$ is a (non-negative) weight.
 Notice that the partition function can also be written as $\mathcal{Z}=\sum_{i}p_{i}$.
@@
 As a Markov chain Monte Carlo simulation, thermalization and autocorrelation effects have to be taken into account to estimate averages and standard errors.
+In the following, we denote by $R$ the number of random METTS states sampled.
+In our calculations, we choose $R=100$ due to our computational budget, from which we obtain error estimates by analyzing thermalization and autocorrelation times as in standard error analysis of Markov chain Monte Carlo methods.
 The METTS algorithm is particularly well-suited to tensor network simulations describing states of low entanglement.
@@
 III.2 Dynamical METTS algorithm in real-time
-Figure 3: Exemplary METTS samples (blue lines) of (a) the correlation functions $\mathcal{C}_{AB}^{i}(t)$ and (b) the spectral functions $\mathcal{C}_{AB}^{i}(\omega)$ of the Shastry-Sutherland model on a $16\times 4$ cylinder for $J/J_{D}=0.63$ and $T/J_{D}=0.1$.
-We investigate the spin correlations and structure factors with $A^{\dagger}=B=S^{z}(\bm{q})$ as in Eqs.
-3 and 4 with $\bm{q}=(\pi,0)$.
-The estimated averages $\mathcal{C}_{AB}(t)$ and $\mathcal{C}_{AB}(\omega)$ with a statistical error estimate from $R=100$ samples are shown as the red line with the shaded light red region.
+Figure 2: Illustration of the dynamical METTS algorithm.
+A sequence of product states $|\sigma_{i}\rangle$ and METTS $|\psi_{i}\rangle$ is sampled according to the thermal METTS algorithm.
+For every METTS the time-dependent correlation function $\mathcal{C}_{AB}^{i}(t)$ as in Eq.
+21 is computed.
+Product states $|\sigma_{i+1}\rangle$ are sampled with transition probability $T_{i\rightarrow i+1}=|\langle\psi_{i}|\sigma_{i+1}\rangle|$.
+From the individual METTS correlation functions $\mathcal{C}_{AB}^{i}(t)$ the spectral function $\mathcal{C}_{AB}(\omega)$ is computed via a Fourier transform and subsequent averaging.
 To study dynamical structure factors we are interested in the conventional time-dependent correlation function $\mathcal{C}_{AB}(t)$ given by,
@@
 $A^{\dagger}=B=S^{z}(\bm{q}),$ (15)
-as defined in Eqs. 3 and 4. A dynamical spectral function is then given as the Fourier transform,
+as defined in Eqs. 4 and 5. A dynamical spectral function is then given as the Fourier transform,
 […]
@@
 $\quad\mathcal{C}_{AB}(\omega)=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\mathcal{C}_{AB}^{i}(\omega),$ (24)
-where,
-[…]
-This suggests the following algorithm to compute dynamical correlation functions $\mathcal{C}_{AB}^{i}(t)$ and $\mathcal{C}_{AB}^{i}(\omega)$.
-We sample METTS states |$\ket{\psi_{i}}$ by sampling from the Markov chain with transition probability $T_{i\rightarrow j}=|\langle\psi_{i}|\sigma_{j}\rangle|^{2}$, just as in the thermal METTS algorithm.
-For every METTS state |$\ket{\psi_{i}}$ we compute $|v_{i}(0)\rangle=B|\psi_{i}\rangle$ and set $|w_{i}(0)\rangle=|\psi_{i}\rangle$ at the beginning.
+{algorithm}
 Choose a random initial product state $|\sigma_{1}\rangle=|\sigma_{1}^{1}\rangle\cdot\ldots|\sigma_{1}^{N}\rangle$
 for $i=1,\ldots,R$ do
-     Compute the METTS state $|\psi_{i}\rangle=\frac{1}{\sqrt{p_{i}}}e^{-\beta H/2}|\sigma_{i}\rangle$
+     1 Compute the METTS state $|\psi_{i}\rangle=\frac{1}{\sqrt{p_{i}}}e^{-\beta H/2}|\sigma_{i}\rangle$
      Compute $|v_{i}(0)\rangle=B|\psi_{i}\rangle$ and set $|w_{i}(0)\rangle=|\psi_{i}\rangle$
@@
 end for
-Algorithm 1 Dynamical METTS for correlators $\mathcal{C}(t)$
+Dynamical METTS for correlators $\mathcal{C}(t)$
+where,
+[…]
+This suggests the following algorithm to compute dynamical correlation functions $\mathcal{C}_{AB}^{i}(t)$ and $\mathcal{C}_{AB}^{i}(\omega)$.
+We sample METTS states |$\ket{\psi_{i}}$ by sampling from the Markov chain with transition probability $T_{i\rightarrow j}=|\langle\psi_{i}|\sigma_{j}\rangle|^{2}$, just as in the thermal METTS algorithm.
+For every METTS state |$\ket{\psi_{i}}$ we compute $|v_{i}(0)\rangle=B|\psi_{i}\rangle$ and set $|w_{i}(0)\rangle=|\psi_{i}\rangle$ at the beginning.
 The next step is to perform a real-time evolution to compute $|v_{i}(t)\rangle$ and $|w_{i}(t)\rangle$ as in Eqs.
-17 and 18.
+18 and 19.
 Notice that the METTS states $|\psi_{i}\rangle$ are not eigenstates of the Hamiltonian $H$, such that the time-evolved states $|v_{i}(t)\rangle$ and $|w_{i}(t)\rangle$ have to be explicitly computed.
@@
 Measurements of $\mathcal{C}_{AB}^{i}(t)$ are only performed at discrete time steps which are a multiple of an elementary step size $\delta_{t}$.
+Figure 3: Exemplary METTS samples (blue lines) of (a) the correlation functions $\mathcal{C}_{AB}^{i}(t)$ and (b) the spectral functions $\mathcal{C}_{AB}^{i}(\omega)$ of the Shastry-Sutherland model on a $16\times 4$ cylinder for $J/J_{D}=0.63$ and $T/J_{D}=0.1$.
+We investigate the spin correlations and structure factors with $A^{\dagger}=B=S^{z}(\bm{q})$ as in Eqs.
+4 and 5 with $\bm{q}=(\pi,0)$.
+The estimated averages $\mathcal{C}_{AB}(t)$ and $\mathcal{C}_{AB}(\omega)$ with a statistical error estimate from $R=100$ samples are shown as the red line with the shaded light red region.
 For convenience, we introduce the shorthand notations,
@@
 $C^{i}(t)=C^{i}_{AB}(t),\quad$ $C^{i}(\omega)=C^{i}_{AB}(\omega).$ (27)
-The algorithm laid out above is summarized in Alg.
-1, and illustrated graphically in Fig.
+The algorithm laid out above is summarized in Fig.
+2, and illustrated graphically in Fig.
 2.
-Using the Shastry-Sutherland model as a concrete example, we show the corresponding correlation function $\mathcal{C}_{AB}^{i}(t)$ and $\mathcal{C}_{AB}^{i}(\omega)$ for several sampled METTS configurations as well as their average value in Fig.
+Using the Shastry-Sutherland model as a concrete example, we show the corresponding correlation function $\mathcal{C}_{AB}^{i}(t)$ and $\mathcal{C}_{AB}^{i}(\omega)$ for several sampled METTS states as well as their average value in Fig.
 3.
+At the given temperature $T/J_{D}=0.1$ we observe a “bunching” of correlation functions $\mathcal{C}_{AB}^{i}(t)$ around the $T=0$ correlation function whereas several other random realizations disperse, illustrating the effect of thermal fluctuations.
 The real-time approach of measuring spectral functions at finite temperature is rigorous as in the limit of long time domain and sufficient sampling of METTS states |$\ket{\psi_{i}}$.
@@
 6 below, the bipartite entanglement entropy $S_{\text{vN}}$ of time evolution based on METTS states of the Shastry-Sutherland model at $T/J_{D}=0.2$ grows much more quickly on average than the one at $T/J_{D}=0.1$.
-Meanwhile, pronounced fluctuations of the entropy after time evolution exist among various sampled METTS configurations.
+Meanwhile, pronounced fluctuations of the entropy after time evolution exist among various sampled METTS states.
 In the limit of infinitely long simulation time $\Omega\rightarrow\infty$ and infinitely accurate calculations of the entangled states |$\ket{\psi_{i}}$, |$\ket{v_{i}(t)}$, and |$\ket{w_{i}(t)}$, this algorithm is exact assuming ergodicity of the METTS Markov chain.
@@
 Our implementation is based on the ITensor library [71] and the TDVP package of ITensor [72].
-Figure 4: Comparison of results from dynamical METTS to the spectral function from exact diagonalization on a $4\times 4$ lattice with fully periodic boundary conditions at $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$ for temperatures $T/J_{D}=0.1$ (a) and $T/J_{D}=0.2$ (b).
-As the exact spectral function from ED is a sum of delta functions, we show Gaussian broadenings for widths $\eta=0.02,0.05,0.10$.
-For the dynamical METTS, we used $R=100$ samples with a real-time evolution up to $\Omega/J_{D}=50$.
-We observe close agreement between the METTS results and the ED results at Gaussian broadening $\eta=0.05$.
 III.4 Spectral analysis and windowing
@@
 III.5 Validation by Exact Diagonalization and DMRG
-Figure 5: Dynamical spin structure factor $S^{z}(\bm{q},\omega)$ of the Shastry-Sutherland model at $J/J_{D}=0.63$ on a $16\times 4$ cylinder obtained using the real-time dynamical METTS algorithm at various temperatures in units of the triplon gap $\Delta=0.458J_{D}$.
-We compare with $S^{z}(\bm{q},\omega)$ at $T=0$ obtained using DMRG for momenta $\bm{q}=(\pi,0)$ (a) and $\bm{q}=(\pi/2,\pi/2)$ (b).
-Shaded regions depict the standard error of the estimator obtained from $R=100$ METTS states with time evolution up to time $\Omega/J_{D}=50$.
-MPS time-evolution has been performed using the 2TDVP algorithm with $\chi_{\text{max}}=2000$.
-We observe that even small features, as well as secondary and tertiary peaks, are well-resolved and converge towards the DMRG result for $T\rightarrow 0$.
+Figure 4: Comparison of results from dynamical METTS to the spectral function from exact diagonalization on a $4\times 4$ lattice with fully periodic boundary conditions at $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$ for temperatures $T/J_{D}=0.1$ (a) and $T/J_{D}=0.2$ (b).
+As the exact spectral function from ED is a sum of delta functions, we show Gaussian broadenings for widths $\eta=0.02,0.05,0.10$.
+For the dynamical METTS, we used $R=100$ samples with a real-time evolution up to $\Omega/J_{D}=50$.
+We observe close agreement between the METTS results and the ED results at Gaussian broadening $\eta=0.05$.
 To demonstrate the accuracy of our method we start by comparing to data on a $4\times 4$ simulation cluster with periodic boundary conditions, where exact diagonalization (ED) can be straightforwardly performed.
@@
 Hence, this benchmark also serves as a good first test of the quality of time-evolution algorithms employed to compute the METTS states as well as the real-time evolution.
+We apply the MPO representing $S^{z}(\bm{q})$ directly for a specific $q$ in our simulations instead of measuring correlations between pairs of sites.
+This is done in order to minimize the amount of measurements needed, and to allow for statistical self-averaging across the lattice during sampling.
 To compute the Fourier transform $S(\bm{q},\omega)$ to frequency space, we compared different windowing functions and found the Hann window Eq.
@@
 Sec.
-III.4.
+III.4 and appendix C.
 For ED, we compare three different values of the broadening $\eta=0.02,0.05,0.10$.
@@
 While both the Gaussian broadening $\eta$ for the ED data as well as the finite time horizon $\Omega$ have a broadening effect on the exact spectrum Eq.
-33, we observe that even small intricate structures of the spectrum (likely stemming from the small system size) are accurately captured by the METTS algorithm.
+34, we observe that even small intricate structures of the spectrum (likely stemming from the small system size) are accurately captured by the METTS algorithm.
 In particular, the positions of the numerous side peaks are accurately reproduced, as well as the overall spectral amplitude.
@@
 Hence, we conclude that having a finite time horizon $\Omega$ as well as only a discrete set of measurements with time steps $\delta_{t}$ only leads to a small broadening of the spectrum.
-Moreover, we report that averaging over only $R=100$ METTS yields a satisfyingly small statistical error indicated by the red-shaded region.
+Discrepancies around the $\omega\rightarrow 0$ limit stem from the finite simulation time-horizon $\Omega$.
+Moreover, we report that averaging over only $R=100$ METTS yields a satisfyingly small statistical error indicated by the shaded regions.
+Figure 5: Dynamical spin structure factor $S^{z}(\bm{q},\omega)$ of the Shastry-Sutherland model at $J/J_{D}=0.63$ on a $16\times 4$ cylinder obtained using the real-time dynamical METTS algorithm at various temperatures in units of the triplon gap $\Delta=0.458J_{D}$.
+We compare with $S^{z}(\bm{q},\omega)$ at $T=0$ obtained using DMRG for momenta $\bm{q}=(\pi,0)$ (a) and $\bm{q}=(\pi/2,\pi/2)$ (b).
+Shaded regions depict the standard error of the estimator obtained from $R=100$ METTS states with time evolution up to time $\Omega/J_{D}=50$.
+MPS time-evolution has been performed using the 2TDVP algorithm with $\chi_{\text{max}}=2000$.
+We observe that even small features, as well as secondary and tertiary peaks, are well-resolved and converge towards the DMRG result for $T\rightarrow 0$.
 We now compare the results from simulations at nonzero temperature with ground state spectral functions from DMRG.
@@
 Our maximal time horizon for the real-time evolution has been chosen as $\Omega/J_{D}=50$.
-Moreover, we collected $R=100$ independent METTS samples for each temperature and observed that this allows sharp error estimates with a large signal-to-noise ratio, as apparent by the small error bars in Fig.
+Moreover, we collected $R=100$ independent METTS states for each temperature and observed that this allows sharp error estimates with a large signal-to-noise ratio, as apparent by the small error bars in Fig.
 5.
@@
 Real-time evolution of MPS is computationally expensive due to unbounded entanglement growth for generic wave functions.
-Imaginary-time correlation functions, on the other hand, can be efficiently computed since for imaginary-time $\tau\rightarrow\infty$ every wavefunction $|\phi\rangle$ will converge to the ground state $|\phi\rangle=|\psi_{0}\rangle$ whenever both wave functions have non-vanishing overlap $|\langle\phi|\psi_{0}\rangle|\neq 0$, i.e.
+Imaginary-time correlation functions, on the other hand, can be efficiently computed since for imaginary-time $\tau\rightarrow\infty$ every wavefunction $|\phi\rangle$ will converge to the ground state $|\phi\rangle=|\psi_{0}\rangle$ whenever both wave functions have non-vanishing overlap $|\langle\phi|\psi_{0}\rangle|\neq 0$, i.e.,
 $\lim\limits_{\tau\rightarrow\infty}e^{-\tau H}|\phi\rangle\simeq|\psi_{0}\rangle,$ (38)
@@
 $\mathcal{U}(z)^{-1}\neq\mathcal{U}(z)^{\dagger}.$ (42)
-There are two conceivable alternatives when defining complex time correlation functions,
+There are two natural alternatives when defining complex time correlation functions,
 $\mathcal{A}_{AB}(z)$ $=\braket{\mathcal{U}(z)^{-1}A\mathcal{U}(z)B}_{\beta}=\braket{e^{iHz}Ae^{-iHz}B}_{\beta},$ (43)
@@
 The Hermitian correlator $\mathcal{H}_{AB}(z)$ as in Eq.
-43, on the other hand, has only recently been investigated in the context of MPS and impurity models [53] and has several intriguing properties which are discussed Sec.
+44, on the other hand, has only recently been investigated in the context of MPS and impurity models [53] and has several intriguing properties which are discussed Sec.
 IV.4.
 We choose the name Hermitian, since whenever $A$ is a Hermitian operator, i.e., $A^{\dagger}=A$, the time-evolved operator $\mathcal{U}(z)^{\dagger}A\mathcal{U}(z)$ is also Hermitian.
-There are computational advantages in evaluating this time-dependent correlation function.
-For convenience, we again introduce the shorthand notations,
+This property is shared with the real-time correlation function, but not with the analytic correlation function, where in general $\mathcal{A}_{AB}(z)^{\star}\neq\mathcal{A}_{BA}(z)$.
+There are computational advantages in evaluating this time-dependent correlation function. For convenience, we again introduce the shorthand notations,
 $\mathcal{A}(z)=\mathcal{A}_{AB}(z),$ (46)
 $\mathcal{H}(z)=\mathcal{H}_{AB}(z).$ (47)
-Figure 7: Sketch of the complex contour on which the analytical correlation function $\mathcal{A}(z)$ is evaluated.
-The maximal imaginary time is chosen to be $\Omega_{\tau}=\beta/2$ in order to allow for an efficient and stable numerical algorithm.
+Figure 7: Sketch of the complex contours on which the analytical correlation function $\mathcal{A}(z)$ and the Hermitian correlation functions $\mathcal{H}(z)$ are evaluated.
+For the analytical correlation function, the maximal imaginary time is chosen to be $\Omega_{\tau}=\beta/2$ in order to allow for an efficient and stable numerical algorithm.
 IV.2 Analytic complex-time correlation functions
 We propose an efficient method for evaluating the analytic complex time correlation function $\mathcal{A}(z)$. Its definition as in Eq. 43 is expressed as,
-$\mathcal{A}(z)$ $=\braket{e^{iHz}Ae^{-iHz}B}_{\beta}$ (47)
+$\mathcal{A}(z)$ $=\braket{e^{iHz}Ae^{-iHz}B}_{\beta}$
 $=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\langle\psi_{i}|e^{izH}Ae^{-izH}B|\psi_{i}\rangle$ (48)
@@
 A naive approach to evaluating $\mathcal{A}(z)$ would be to compute intermediate states,
-$|v_{i}(t)\rangle=e^{-iHz}B|\psi_{i}\rangle=e^{-iHt}e^{-\tau H}B|\psi_{i}\rangle$ (55)
-$|w_{i}(t)\rangle=e^{-iH\overline{z}}|\psi_{i}\rangle=e^{-iHt}e^{+\tau H}|\psi_{i}\rangle,$ (56)
-with $z=t-i\tau$ similar to the real-time correlation algorithm Alg.
-1.
+$|v_{i}(z)\rangle=e^{-iHz}B|\psi_{i}\rangle=e^{-iHt}e^{-\tau H}B|\psi_{i}\rangle$ (55)
+$|w_{i}(z)\rangle=e^{-iH\overline{z}}|\psi_{i}\rangle=e^{-iHt}e^{+\tau H}|\psi_{i}\rangle,$ (56)
+with $z=t-i\tau$ similar to the real-time correlation algorithm Fig.
+2.
 However, this approach is plagued by severe numerical instability when computing the state $|w_{i}(t)\rangle$.
 Computing a positive exponential $e^{+\tau H}|\psi_{i}\rangle$ is numerically highly unstable and prohibits most practical calculations.
+This is due to the fact that the initial METTS state is a low-energy state.
+High-energy contributions are truncated during the time evolution but would need to be resolved accurately when performing positive time evolution causing severe numerical instabilities.
 When restricting our time horizon to $\Omega_{\tau}=\beta/2$ we can circumvent ever having to compute positive real exponential in the context of the METTS algorithm.
+Thus, given we are only interested in simulating complex times with $\Omega_{\tau}<\beta/2$ we propose that the analytic correlator can be evaluated in a numerically stable way.
+This is done by absorbing part of the positive time evolution into the negative time evolution performed for the METTS states.
+Notice that a smaller complex angle $\theta$ allows a longer real-time component $\Omega_{t}=\Omega\cos\theta$ while still fulfilling $\Omega_{\tau}=\Omega\sin\theta<\beta/2$.
 The key idea is that the positive imaginary time evolution can be absorbed by the negative imaginary time evolution performed when calculating an individual METTS |$\ket{\psi_{i}}$.
-For this, we rewrite Eq.
-47 as,
-Algorithm 2 Dynamical METTS for analytic correlator $\mathcal{A}(z)$
+For this, we rewrite Sec.
+IV.2 as,
+$\mathcal{A}$ $(z)=\braket{e^{izH}Ae^{-izH}B}_{\beta}=$
+$=\braket{e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}}_{\beta}$
+$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle$
+$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\;\mathcal{A}^{i}(z),$ (57)
+where we defined the sample analytic correlation function as,
+$\mathcal{A}^{i}(z)\equiv\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle.$ (58)
+Introducing the factor $e^{\pm i\Omega_{t}H}$ as in Eq. 58 allows to rewrite $\mathcal{A}^{i}(z)$ in a simple form,
+$\mathcal{A}^{i}(z)$ $=\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle$
+$=\frac{1}{p_{i}}\langle\sigma_{i}|e^{-i(\Omega_{z}-z)H}Ae^{-izH}Be^{i\overline{\Omega}_{z}H}|\sigma_{i}\rangle.$ (59)
+For $k=0,\ldots,S$ we introduce the auxiliary computational states |$\ket{\phi_{i}^{k}}$ and |$\ket{\chi_{i}^{k}}$ defined as,
+|$\ket{\phi_{i}^{k}}$ $\equiv\frac{1}{\sqrt{p_{i}}}e^{ik\overline{\delta}_{z}H}\ket{\sigma_{i}}$ $=\frac{1}{\sqrt{p_{i}}}e^{ik{\delta_{t}}H}e^{-k{\delta_{\tau}}H}\ket{\sigma_{i}},$ (60)
+|$\ket{\chi_{i}^{k}}$ $\equiv e^{-ik\delta_{z}H}B\ket{\phi_{i}^{S}}$ $=e^{-ik\delta_{t}H}e^{-k\delta_{\tau}H}B\ket{\phi_{i}^{S}}.$ (61)
+Importantly, we see that these states can all be computed by performing a negative imaginary time evolution, i.e., computing $e^{-\tau H}\ket{\psi}$ for some state |$\ket{\psi}$.
+This insight is the main reason why $\mathcal{A}^{i}(z)$ can be computed efficiently.
+Notice that,
+$\bra{\phi_{i}^{S-k}}$ | $=\frac{1}{\sqrt{p_{i}}}\bra{\sigma_{i}}e^{-i(\Omega_{z}-k\delta_{z})H},$ (62)
+|$\ket{\phi_{i}^{S}}$ $=\frac{1}{\sqrt{p_{i}}}e^{i\overline{\Omega}_{z}H}|\sigma_{i}\rangle.$ (63)
+Using a discretized $z=k\delta_{z}$ we can rewrite Eq. 59 as,
+$\mathcal{A}^{i}(k\delta_{z})=\bra{\phi^{S-k}_{i}}A\ket{\chi_{i}^{k}}.$ (64)
+{algorithm}
+[t!] Dynamical METTS for analytic correlator $\mathcal{A}(z)$
 1:Choose a random initial product state $|\sigma_{1}\rangle=|\sigma_{1}^{1}\rangle\cdot\ldots|\sigma_{1}^{N}\rangle$
@@
 19:end for
-$\mathcal{A}$ $(z)=\braket{e^{izH}Ae^{-izH}B}_{\beta}=$ (57)
-$=\braket{e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}}_{\beta}$ (58)
-$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle$ (59)
-$=\frac{1}{\mathcal{Z}}\sum_{i}p_{i}\;\mathcal{A}^{i}(z),$ (60)
-where we defined the sample analytic correlation function as,
-$\mathcal{A}^{i}(z)\equiv\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle.$ (61)
-Introducing the factor $e^{\pm i\Omega_{t}H}$ as in Eq. 61 allows to rewrite $\mathcal{A}^{i}(z)$ in a simple form,
-$\mathcal{A}^{i}(z)$ $=\langle\psi_{i}|e^{-i\Omega_{t}H}e^{izH}Ae^{-izH}Be^{i\Omega_{t}H}|\psi_{i}\rangle$ (62)
-$=\frac{1}{p_{i}}\langle\sigma_{i}|e^{-i(\Omega_{z}-z)H}Ae^{-izH}Be^{i\overline{\Omega}_{z}H}|\sigma_{i}\rangle.$ (63)
-For $k=0,\ldots,S$ we introduce the auxiliary computational states |$\ket{\phi_{i}^{k}}$ and |$\ket{\chi_{i}^{k}}$ defined as,
-|$\ket{\phi_{i}^{k}}$ $\equiv\frac{1}{\sqrt{p_{i}}}e^{ik\overline{\delta}_{z}H}\ket{\sigma_{i}}$ $=\frac{1}{\sqrt{p_{i}}}e^{ik{\delta_{t}}H}e^{-k{\delta_{\tau}}H}\ket{\sigma_{i}},$ (64)
-|$\ket{\chi_{i}^{k}}$ $\equiv e^{-ik\delta_{z}H}B\ket{\phi_{i}^{S}}$ $=e^{-ik\delta_{t}H}e^{-k\delta_{\tau}H}B\ket{\phi_{i}^{S}}.$ (65)
-Importantly, we see that these states can all be computed by performing a negative imaginary time evolution, i.e., computing $e^{-\tau H}\ket{\psi}$ for some state |$\ket{\psi}$.
-This insight is the main reason why $\mathcal{A}^{i}(z)$ can be computed efficiently.
-Notice that,
-$\bra{\phi_{i}^{S-k}}$ | $=\frac{1}{\sqrt{p_{i}}}\bra{\sigma_{i}}e^{-i(\Omega_{z}-k\delta_{z})H},$ (66)
-|$\ket{\phi_{i}^{S}}$ $=\frac{1}{\sqrt{p_{i}}}e^{i\overline{\Omega}_{z}H}|\sigma_{i}\rangle.$ (67)
-Using a discretized $z=k\delta_{z}$ we can rewrite Eq. 63 as,
-$\mathcal{A}^{i}(k\delta_{z})=\bra{\phi^{S-k}_{i}}A\ket{\chi_{i}^{k}}.$ (68)
 To evaluate these expectation values for all discretized values of $z=k\delta_{z}$ ($k=0,\ldots,S$), we compute the states |$\ket{\phi_{i}^{k}}$ for $k=0,\ldots,S$ successively by complex time evolution and store all intermediate states to disk for later use.
@@
 $\mathcal{A}(z)$ is then obtained by averaging over $\mathcal{A}^{i}(z)$.
-We summarize the described algorithm in Alg.
+We summarize the described algorithm in Sec.
+IV.2.
+Sec.
+IV.2 is in several aspects similar to Fig.
 2.
-Alg.
-2 is in several aspects similar to Alg.
-1.
 Both algorithms require computing the METTS |$\ket{\psi_{i}}$ and then time-evolves two further states to compute time-dependent correlators.
 Moreover, the basic thermal METTS Markov chain is constructed in both cases.
-A difference in Alg.
-2 is the necessity of storing the auxiliary states |$\ket{\phi_{i}^{k}}$.
+A difference in Sec.
+IV.2 is the necessity of storing the auxiliary states |$\ket{\phi_{i}^{k}}$.
 On present-day computers, this typically neither imposes severe time nor memory constraints.
 However, in the limit of small angle $\theta$ the memory requirement can be considerable due to the large number of discrete time steps and requires corresponding storage capacities.
+Figure 8: Average bipartite von Neumann entanglement entropy $S_{\text{vN}}$ of the auxiliary state |$\ket{\phi_{i}^{k}}$ Eq.
+60 for evaluating the analytical correlation function $\mathcal{A}(z)$.
+We consider the $16\times 4$ Shastry-Sutherland lattice for different values of the complex angle $\theta$ and $J/J_{D}=0.63$ and compare results for temperatures (a) $T/J_{D}=0.1$ and (b) $T/J_{D}=0.2$.
+The average is computed from $R=100$ METTS states and the shaded region shows the estimated standard deviation of the mean.
 The key quantity determining the efficiency of the algorithm is the bipartite von-Neumann entanglement entropy $S_{\text{vN}}$ as in Eq.
@@
 Hence, the maximal complex times $|z|$ differ for different values of $\theta$.
-Figure 8: Average bipartite von Neumann entanglement entropy $S_{\text{vN}}$ of the auxiliary state |$\ket{\phi_{i}^{k}}$ Eq.
-64 for evaluating the analytical correlation function $\mathcal{A}(z)$.
-We consider the $16\times 4$ Shastry-Sutherland lattice for different values of the complex angle $\theta$ and $J/J_{D}=0.63$ and compare results for temperatures (a) $T/J_{D}=0.1$ and (b) $T/J_{D}=0.2$.
-The average is computed from $R=100$ METTS samples and the shaded region shows the estimated standard deviation of the mean.
+In our simulations, the maximal computed entanglement entropy exhibits a plateau due to simulating at finite bond dimension.
 Figure 9: Same as Fig. 8, but for the auxiliary state |$\ket{\chi_{i}^{k}}$ Eq. 61.
 IV.3 Stochastic analytic continuation
-Figure 10: Dynamical structure factors ${S}(\bm{q},\omega)$ from stochastic analytic continuation based on analytic correlation functions $S(\bm{q},z)$ at several complex angles.
-Here the spectrum in obtained from sampling at $\alpha=\alpha_{c}$.
-Data is shown for the $16\times 4$ lattice with $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$.
-We compare between temperatures $T/J_{D}=0.1$ (a) and $T/J_{D}=0.2$ (b).
-Figure 11: Same as Fig. 10, where the spectrum is obtained via weighted averaging the collected histogram from $\log\alpha=0$ to $\log\alpha=-6$.
 To retrieve the spectral function from $\mathcal{A}(z)$ we perform analytic continuation, for which various approaches have been proposed and tested in the context of quantum Monte Carlo simulations [74, 75, 76, 77].
@@
 Generally, inverse Laplacian approaches from imaginary time correlation functions lead to a biased result where the details of the efficiency are case-dependent.
-In most cases, dominant peaks capturing well-defined quasiparticles can be resolved in a satisfactory way.
-However, for a spectrum with contributions from both a low energy peak and a higher energy continuum, numerical continuation is hard to quantitatively resolve the information at high frequency.
-On the other hand, we propose the numerical analytic continuation to be more efficient via correlation functions defined on the complex time case, since real-time contributions of the correlator, even for short periods of time, carry information about the high-frequency spectrum.
+In most cases, dominant peaks capturing well-defined quasiparticles can be resolved satisfactorily.
+However, for a spectrum with contributions from both a low energy peak and a higher energy continuum, numerical continuation faces difficulties in quantitatively resolving the information at high frequency.
+On the other hand, we propose the numerical analytic continuation to be more efficient via correlation functions defined on the complex time case.
+At first sight, this natural extension would improve the performance of conventional imaginary-time-based numerical analytic continuation approaches since Green’s functions that we use as input are closer to the desired real frequency angle.
+Most importantly, the imaginary time Green’s function is dominated by the low-energy spectrum of the system.
+Instead, the high-energy part of the spectrum contributes significantly to the real-time correlator.
+Hence analytic continuations in our general consideration, with real-time contributions of the correlator, even for short periods, may give an improvement in resolving the spectrum.
 We now document our approach to stochastic analytic continuation (SAC) based on correlation functions in the complex time plane.
@@
 $P(A)\propto e^{-\frac{\chi^{2}(A)}{\alpha}},$ (70)
-where $\alpha$ denotes the fictitious inverse temperate, and $\chi^{2}(A)$ is the $\chi^{2}$ that measures the deviation of correlation function based on any given $A(\omega)$ during stochastic sampling, with the physical one that is observed from METTS calculation.
+where $\alpha$ denotes the fictitious temperature, and $\chi^{2}(A)$ is the $\chi^{2}$ that measures the deviation of correlation function based on any given $A(\omega)$ during stochastic sampling, with the physical one that is observed from METTS calculation.
 In the current case, fluctuations from both real and imaginary parts of the correlation function are included in the definition of $\chi^{2}$,
@@
 $\sum_{i=1}^{R}\text{Im}(\mathcal{A}^{i}(z_{k})-\overline{\mathcal{A}}(z_{k}))\text{Im}(\mathcal{A}^{i}(z_{k^{\prime}})-\overline{\mathcal{A}}(z_{k^{\prime}})),$
-where $\mathcal{A}^{i}(z_{k})$ is the correlation function measured from the $i$-th METTS configuration.
+where $\mathcal{A}^{i}(z_{k})$ is the correlation function measured from the $i$-th METTS state.
 So far, the major difference between the current setup and the standard analytic continuation approach based on imaginary time correlation functions is the complexity of time $z$, correlation function $G(z)$, and the integration kernel $K(z,\omega)$.
@@
 [76].
+Figure 10: Dynamical structure factors ${S}(\bm{q},\omega)$ from stochastic analytic continuation based on analytic correlation functions $S(\bm{q},z)$ at several complex angles.
+Here the spectrum in obtained from sampling at $\alpha=\alpha_{c}$.
+Data is shown for the $16\times 4$ lattice with $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$.
+We compare between temperatures $T/J_{D}=0.1$ (a) and $T/J_{D}=0.2$ (b).
+Figure 11: Same as Fig. 10, where the spectrum is obtained via weighted averaging the collected histogram from $\log\alpha=0$ to $\log\alpha=-6$.
 We show results of analytic continuation at $T/J_{D}=0.1$ and $0.2$ for several complex angles in Figs.
@@
 First, the spectrum obtained via the sampling at $\alpha_{c}$ does not differ from the one via weighted averaging the data from high to low temperature.
-Clearly, the position of the lowest triplon mode at $T/J_{D}=0.1$ is constant with both perturbation theory and the one from hermitian correlation function (at $\theta\longrightarrow 0$ limit) as shown in Fig.
+Clearly, the position of the lowest triplon mode at $T/J_{D}=0.1$ is constant with both perturbation theory and the one from the Hermitian correlation function (at $\theta\longrightarrow 0$ limit) as shown in Fig.
 12.
@@
 On the other hand, we did not observe a significant $\theta$ dependence for the thermal broadened spectrum at $T/J_{D}=0.2$.
+The optimal angle for calculating the analytic correlation function is a question of balancing the increased computational cost at a smaller complex angle and the higher uncertainty in analytical continuation at a larger one.
+Certainly, among the numerical cost-accessible regions in the complex time plane, one may push the calculation to the smallest possible complex angle to acquire the most accurate spectral function.
 IV.4 Hermitian complex-time correlation functions
@@
 We now focus our attention on the second possible generalization of correlation functions to complex time, the Hermitian correlation function as defined in Eq.
-43, which expanded in the METTS basis yields,
+44, which expanded in the METTS basis yields,
 $\mathcal{H}_{AB}(z)$ $=\braket{e^{iH\overline{z}}Ae^{-iHz}B}$ (74)
@@
 $\mathcal{H}^{i}_{AB}(z)\equiv\langle\psi_{i}|e^{iH\overline{z}}Ae^{-iHz}B|\psi_{i}\rangle.$ (77)
-To evaluate this correlation function we can proceed analogously to the conventional time-dependent correlation function algorithm in Alg. 1, and define the states,
+To evaluate this correlation function we can proceed analogously to the conventional time-dependent correlation function algorithm in Fig. 2, and define the states,
 $|v_{i}(z)\rangle=e^{-iHz}B|\psi_{i}\rangle=e^{-H\tau}e^{-iHt}B|\psi_{i}\rangle,$ (78)
@@
 $\mathcal{H}^{i}_{AB}(z)=\langle w_{i}(z)|A|v_{i}(z)\rangle,$ (80)
-using Alg.
+using Fig.
 2 adjusted for the complex time evolution along the contour $z=|z|e^{-i\theta}$ shown in Fig.
 7.
-A key difference to the dynamical METTS algorithm Alg.
+A key difference to the dynamical METTS algorithm Fig.
 2 is the fact that the time evolution operators $e^{-iHz}$ and $e^{iH\overline{z}}$ are not norm preserving.
@@
 12.
-Remarkably, for $T/J_{D}=0.1$ the position of the dominant peak at $\omega/J_{D}\approx 0.46$ is always obtained consistently even for sizable values of the complex angle $\theta$.
+Remarkably, for $T/J_{D}=0.1$ the position of the dominant peak at $\omega/J_{D}\approx 0.458$ is always obtained consistently even for sizable values of the complex angle $\theta$.
 Meanwhile, the height and width of the dominant peak are well captured as long as $\theta\leq\pi/16$.
@@
 12(b) is generally well captured at larger values of $\theta$.
-However, the peak around the triplon gap for $\omega/J_{D}\approx 0.46$ is overestimated at finite angles of $\theta$.
+However, the peak around the triplon gap for $\omega/J_{D}\approx 0.458$ is overestimated at finite angles of $\theta$.
 Figure 13: Average bipartite von Neumann entanglement entropy $S_{\text{vN}}$ of the states |$\ket{\tilde{v}_{i}(z)}$ as in Eq.
 84 for computing the Hermitian correlator Eq.
-43.
+44.
 Data on the $16\times 6$ Shastry-Sutherland lattice for different values of the complex angle $\theta$ at $J/J_{D}=0.63$ and $\bm{q}=(\pi/2,\pi/2)$.
 We compare results for temperatures $T/J_{D}=0.1$ (a) and $T/J_{D}=0.2$ (b).
-For larger values of $\theta$ we observe convergence of $S_{\text{vN}}$ towards the ground state entanglement entropy.
+For larger values of $\theta$ we observe convergence of $S_{\text{vN}}$ towards the ground-state entanglement entropy.
 Our main motivation for employing the Hermitian correlation function is to limit the entanglement when time-evolving an MPS state.
@@
 On the other hand, the definition of the Hermitian correlator has the advantage of easy implementation and the imaginary time evolution dissipates the entanglement entropy dramatically.
-Meanwhile, the main structure of the real frequency spectrum is well captured even in the presence of a systemic deviation due to imaginary contributions.
+Meanwhile, the main structure of the real-frequency spectrum is well captured even in the presence of a systemic deviation due to imaginary contributions.
 Only recently, this version of a complex time correlator has been studied [53] in the context of impurity models with MPS.
-It is worth pointing out, that the idea of introducing an imaginary-time component to reduce entanglement is similar in spirit to the dissipation-assisted operator evolution proposed in Ref.
+Moreover, Ref.
+[53] provides a method of extrapolating correlation functions from the complex plane to the real axis which demonstrably is capable of achieving high accuracy.
+It is worth pointing out that the idea of introducing an imaginary-time component to reduce entanglement is similar in spirit to the dissipation-assisted operator evolution proposed in Ref.
 [45].
-The imaginary-time component acts as a natural ”dissipator”.
-One advantage of the present approach is that existing implementations of time-evolution algorithms, such as TDVP, can be readily employed to compute complex-time evolution.
+The imaginary-time component acts as a natural “dissipator”.
+One advantage of the present approach is that existing implementations of time-evolution algorithms, such as TDVP, can be readily employed to compute complex time evolution.
 The SAC method based on the complex plane correlation function has still leftover possibilities to be improved in the future: one promising suggestion is to perform continuation via combined information of correlators at several angles along the complex plane.
-Apart from the stochastic approach, an interesting direction is to combine the Nevanlinna analytic continuation [78] approach with the correlation functions obtained at certain METTS snapshots.
+Apart from the stochastic approach, an interesting direction is to combine the Nevanlinna analytic continuation [78] approach with the correlation functions obtained at certain METTS states.
 On the other hand, a common technique for computing dynamical spectral functions at zero temperatures by using real-time evolution with MPS, is to perform linear prediction in order to extend the time-series of correlation functions [79, 39, 80, 81].
@@
 It would be interesting to find out whether this method would be beneficial in the present context to extend the METTS time series.
+The scaling of the method is determined by the maximal MPS bond dimension $\chi$ which needs to be simulated and the MPO bond dimension $D$ which typically remains finite.
+To perform both the METTS sampling as well as the time-evolution of METTS states, the proposed method has the same scaling as DMRG, which is $\mathcal{O}(\chi^{3}D+\chi^{2}D^{2})$, since we are using the TDVP time evolution method.
+The key novelty is, that due to the complex time evolution, the maximal necessary MPS bond dimension $\chi$ needed to resolve the time-evolved state is bounded, as opposed to real-time evolution, where the entanglement in general grows unbounded.
 In summary, we demonstrated that the combination of METTS with either real or complex-time evolution can be applied to study spectral functions of two-dimensional quantum many-body systems using MPS.
@@
 Figure 14: Imaginary time correlation function using TEBD and the 2-site TDVP approach, for $T/\Delta=0$ (a), $0.1$ (b) and $0.2$ (c).
-Two dashed lines, including $c_{2}e^{-\Delta\tau}$ ( $\Delta\approx 0.459$ ) and $c_{1}e^{-0.56\tau}$ are guides to the eye.
+Two dashed lines, including $c_{2}e^{-\Delta\tau}$ ( $\Delta\approx 0.458$ ) and $c_{1}e^{-0.56\tau}$ are guides to the eye.
 The time evolution approach for measuring correlation function at $T=0$ is based on the ground state via DMRG calculation, where the ones at $T/J_{D}=0.1$ and $0.2$ are from averaging the METTS configurations.
@@
 It turns out to be a serious issue for the Shastry-Sutherland model when the temperature is much lower than the first triplon gap $\Delta$.
-In this low energy case, the MPS during the time evolution step does not differ significantly from the ground state, which is known to be a product state of spin singlets living on the off-diagonal bonds.
+In this low-energy case, the MPS during the time evolution step does not differ significantly from the ground state, which is known to be a product state of spin singlets living on the off-diagonal bonds.
 The projection error leads to a biased estimation of the spin gap, even with the 2-site TDVP approach.
@@
 14.
-At low temperatures, one expects the imaginary time correlation function to decay with an exponential factor based on the estimated spin gap ($\Delta\approx 0.459J_{D}$) from the perturbation theory.
+At low temperatures, one expects the imaginary time correlation function to decay with an exponential factor based on the estimated spin gap ($\Delta\approx 0.458J_{D}$) from the perturbation theory.
 However, as can be seen from the plot, $G(\tau)$ obtained from the 2-site TDVP method deviates from this expectation with a much steeper slope, for $T/\Delta=0$ and $0.1$.
@@
 On the other hand, the overall shape of the spectrum at $\theta=\pi/8$ is much more stable, although we observe an unphysical splitting of the second peak at small $\alpha$.
+Appendix C Comparison of windowing functions
+Figure 16: Comparison of spectral functions obtained via ED and the real-time METTS algorithm as in Fig. 4, but using a Parzen windowing function Eq. 91.
+In this appendix we compare the results of computing the spectral functions using different windowing functions to perform the Fourier transform.
+In the main text in Fig.
+4 we used the Hann window of the form,
+$w(x)=0.5+0.5\cos\left(\pi x\right)\quad\text{for}\quad 0\leq|x|\leq 1.$ (90)
+One could ask to what extent the result depends on the windowing function.
+To answer this question we compare two different windowing functions.
+First, we show results using the Parzen window of the form,
+$w(x)=\begin{cases}1-6|x|^{2}+6|x|^{3}\quad\text{if}\quad 0\leq|x|\leq 1/2\\ 2(1-|x|^{3})\quad\text{if}\quad 1/2\leq|x|\leq 1\end{cases},$ (91)
+in Fig. 16. The Parzen windowing function has previously been used in the context of time-dependent DMRG at zero temperature [20].
+Figure 17: Comparison of spectral functions obtained via ED and the real-time METTS algorithm as in Fig. 4, but using a Gaussian windowing function Eq. 92 of width $\sigma=1/2$.
+An equally valid choice of windowing function is a Gaussian window of the form,
+$w(x)=\exp\left(-\frac{|x|^{2}}{2\sigma^{2}}\right)\quad\text{for}\quad 0\leq|x|\leq 1,$ (92)
+where $\sigma$ controls the width of the Gaussian window. Results from applying the Gaussian windowing function with $\sigma=1/2$ are shown in Fig. 17.
+Overall, we see that due to the chosen maximal simulation time $\Omega/J_{D}=50$ our results are not strongly affected by the choice of windowing function.
+The effects of the windowing function are expected to vanish for $\Omega\rightarrow\infty$.